### v2

# Romance, Relief, and Regret: Teen Narratives of Chatbot Overreliance

Abstract
As Generative Artificial Intelligence (GenAI) driven chatbots like Character.AI become embedded in adolescent life, they raise concerns about emotional dependence and digital overreliance. While studies have investigated the overreliance of adults on these chatbots, they have not investigated teens’ interactions with chatbots with customizable personas. We analyzed 318 Reddit posts made by users self-reported as 13-17 years old on the Character.AI subreddit to understand patterns of overreliance. We found teens commonly begin using chatbots for emotional support or creative expression, but many develop strong attachments that interfere with offline relationships and daily routines. Their posts revealed recurring signs of psychological distress, cycles of relapse, and difficulty disengaging. Teens reported that their overreliance often ended when they reflect on the harm, return to in-person social settings, or become frustrated by platform restrictions. Based on the implications of our findings, we provide recommendations for future chatbot design so they can promote self-awareness, support real-world engagement, and involve teens in developing safer digital tools.
Introduction
Recently, AI companion apps have surged in popularity among teens, with media coverage highlighting how people describe their interactions with chatbots as emotionally meaningful—sometimes even romantic (Murphy 2025). Platforms like Character.AI111https://character.ai are frequently portrayed as offering on-demand conversations with highly responsive virtual characters (Breen 2025). There are other reported perceived benefits of conversational AI, such as reducing loneliness, offering a safe space for self-expression, and providing nonjudgmental support during stressful periods (Xie, Pentina, and Hancock 2023; Salah et al. 2023).
However, these benefits can also drive patterns of digital addiction (referred to in this paper as overreliance), a phenomenon observed in other digital technologies (Lemmens, Valkenburg, and Peter 2009; Griffiths and Kuss 2017). Behavioral addiction is hard to define because everyday digital activities are not inherently harmful, so researchers instead examine factors like disruption to daily life, emotional escape, and loss of control (Lemmens, Valkenburg, and Peter 2009; Griffiths and Kuss 2017). Despite smartphones or social media advantages such as accessibility, convenience, and ability to connect with others, these technologies have also introduced a range of negative consequences—users frequently struggle to regulate their screen time, experience disruptions to sleep and daily routines, and report declines in mental health (Lemmens, Valkenburg, and Peter 2009). Overreliance during sensitive developmental stages can have especially harmful effects, with adolescents identified as a particularly vulnerable group (Griffiths and Kuss 2017).
Teens are especially susceptible to developing problematic relationships with technology due to ongoing cognitive and emotional development (Chen et al. 2022). Prior research has shown that teens often struggle with overreliance on social media, using platforms compulsively to seek validation or cope with stress, sometimes resulting in increased anxiety, depression, and poor self-esteem (Duffy 2025). Similarly, recent studies have identified that users may also have overreliance AI tools like ChatGPT—primarily for academic help or companionship—which in some cases leads to decreased creativity, poor study habits, and emotional detachment from real-life relationships (Zhang et al. 2024; Hu, Mao, and Kim 2023). These studies focus mostly on adults or university students, overlooking teens who may be more emotionally sensitive and developmentally vulnerable. While previous research focused on teens overreliance on social media and gaming (Giordano, Schmit, and McCall 2023), yet, a critical question remains: how are teens disclosing and contextualizing their growing reliance on personalizable conversational AI companions such as Character.AI? What differentiates generative AI companions from other technologies is the parasocial relationships without presence of humans and being highly engaging emotionally responsive human-like conversations (Maeda and Quan-Haase 2024). There has been media coverage and anecdotal reports suggesting emotional attachments and overreliance forming between teens and these AI companions, however, there is little empirical research examining this phenomenon. Therefore, we ask the following research questions:
RQ1: How do teens describe their initial reasons for engaging with Character.AI on Reddit?
RQ2: How do teens’ descriptions of their Character.AI use align with patterns of behavioral addiction?
RQ3: How do teens describe their decisions to disengage from or reduce their use of Character.AI?
To address these questions, we qualitatively analyzed 318 reddit/charcter.AI posts created by users who reported themselves being 13-17 years old. We found that teens initially engaged with Character.AI as a personal coping tool, turning to it not just for entertainment but as a non-judgmental space to manage mental health struggles, or fill emotional voids or a lack of supportive relationships in their offline lives (RQ1). Teens’ reported experiences with Character.AI revealed patterns of their compulsive use, emotional dependence, and distress that often disrupted their daily lives and real-world relationships (RQ2). Teens described disengaging from Character.AI either through internal realizations about its negative impact on their well-being or due to external factors such as reengaging with real-life relationships or frustrations with platform restrictions, often expressing a desire for healthier habits and emotional support (RQ3).
Our study makes the following novel contributions to the AI ethics and human-centered computing research communities: 1) we present one of the first empirical grounding of teen disclosures about AI companion overreliance, mapping their experiences to the six components of behavioral addiction (salience, mood modification, tolerance, withdrawal, conflict, relapse), 2) we identified emotional and relational dynamics unique to teen overreliance on AI companions which involves intense emotional attachment and simulated intimacy, making it distinct from previously studied forms of digital overuse such as gaming or social media, and 3) we provide Design-relevant insights for safer AI platforms by analyzing how teens reflect on disengagement. We surface user-informed strategies for future chatbot design—especially features that promote emotional regulation, real-world social engagement, and usage awareness.
Background
Adolescent Technology Behavioral Addiction
Each new digital platform—social media, gaming, or streaming—brings renewed concerns about its effect on teens (Griffiths and Kuss 2017; Griffiths 2010). Most teens use technology without issues, but some face emotional distress, loneliness, low life satisfaction, social withdrawal, or family conflict due to excessive use (Lemmens, Valkenburg, and Peter 2009). For these individuals, technology stops being just a functional tool or hobby and starts functioning more like emotional companions (Bopp et al. 2019), potentially leading to behavioral addiction. Defining behavioral addiction is difficult because the behaviors involved, such as playing video games or using social media, are not inherently harmful and are often part of daily life, especially for adolescents (Griffiths and Kuss 2011). As a result, researchers focus on how and why the behavior is used—whether it disrupts daily life, helps escape negative emotions, or persists despite loss of control (Lemmens, Valkenburg, and Peter 2009; Griffiths and Kuss 2017).
One widely used framework identified six key features that are usually present in behavioral addiction (Griffiths 2005). First is salience—when the activity becomes the most important thing in a person’s life, dominating their thoughts and daily routine. Second is mood modification—using the activity to change one’s emotional state, like seeking comfort, excitement, or relief from stress. Third is tolerance, which means needing more of the activity over time to get the same emotional effect. Fourth is withdrawal symptoms, which can include anxiety, irritability, or sadness when the activity is reduced or stopped. Fifth is conflict—when the behavior starts damaging relationships, school performance, or personal goals. Finally, there’s relapse, where the individual tries to cut back or quit but eventually returns to the behavior, often with even more intensity. We examined Reddit posts guided by these six signs to understand AI overreliance and distinguish between enthusiastic and problematic use. We actively choose to avoid using the term “addiction” throughout the paper because of the its associated stigmatized notion (Corrigan et al. 2017).
Platform Design, Moderation, and Youth Safety
Many studies have focused on reducing or managing the risks that technology use poses to young people, especially teens. While some harms, like social pressures or mental health impacts, might be difficult to prevent entirely, others are more connected to how technologies are designed. Design patterns help us understand how digital systems function. Although these patterns are usually considered neutral, they can carry ethical weight because how a system is designed can reflect certain values or intentions (Flanagan et al. 2011). Bad design can result from inexperience, but “dark patterns” are intentionally harmful (Gray et al. 2018). These design choices are meant to manipulate users, often leading to negative experiences (Zagal, Björk, and Lewis 2013). Past research has identified these tactics as contributing to tech overuse and even addiction. For instance, random reward systems (e.g., loot boxes), emotional triggers (e.g., likes and comments), or strategies that prevent users from pausing (e.g., autoplay or countdowns) are all tactics that can encourage compulsive use (Flayelle et al. 2023).
Aside from how platforms are designed, researchers have explored ways to regulate and monitor how people use technology. One common approach is self-monitoring, where users reflect on their behavior and try to improve it (Park et al. 2021). This might involve using apps that track screen time or send reminders to take breaks. Although helpful, many users—especially teens—struggle to stay motivated and often give up using them. Parental involvement has since become a major focus in youth online safety efforts. Many digital safety tools are designed with parents in mind (Wisniewski et al. 2017). However, research also shows that many teens feel these tools are too restrictive and do not respect their need for independence (Ghosh et al. 2018). Past work suggests that giving teens more autonomy and involving them in the design process could lead to better outcomes (Wisniewski et al. 2017). Instead of simply limiting what teens can do, newer approaches are trying to support healthy habits and decision-making by encouraging self-awareness and shared responsibility between teens, parents, and platform designers (Chen et al. 2022).
Chatbots and Socio-Emotional Dependency
Advances in AI have made chatbots popular, especially among teens (Sidoti, Park, and Gottfried 2025). Human-like chatbots tend to generate higher user satisfaction and dependency (Yu and Lan 2024). While some users report reduced loneliness, the risks of emotional reliance remain underexplored. Replika’s emotional design has led to deep attachments, even when seen as manipulative or harmful (Laestadius et al. 2024). Teens may be especially vulnerable. Some adolescents acknowledge that using chatbots for schoolwork could harm their critical thinking (Pratt, Madhavan, and Weleff 2024), yet there is limited understanding of when students are developmentally ready for AI in education (Dergaa et al. 2024). Psychological frameworks such as attachment theory help explain why users form emotionally dependent relationships with AI companions (Xie, Pentina, and Hancock 2023). Previous research on parasocial relationships and interactions can also be used to explain the nature of these interactions. Originally used to describe the relationship between a spectator (a viewer or user) and performer (usually a mass media personality), these relationships are one-sided, non-dialect, and not susceptible to mutual development (Horton and Richard Wohl 1956). With the advancements of AI, these type of relationships may become blurred since interactions with chatbots can appear mutual. Understanding these parasocial relationships can also provide insight on the potential loss that users may experience with chatbots (Cohen 2003). People with anxiety or loneliness often see emotionally intelligent chatbots as safe havens, which can lead to rumination and overuse (Hu, Mao, and Kim 2023). Concerns also extend to content safety. Replika has been reported to engage in sexually suggestive exchanges with minors (Namvarpour, Pauwels, and Razi 2025), raising concerns about age-appropriate design (Park, Singh, and Wisniewski 2023). Risks from GenAI include inappropriate advice, exposure to violent or sexual content, misinformation, and privacy threats (Yu et al. 2025). These risks underscore the need for ongoing oversight and study of adolescent chatbot use.
Some scholars caution against labeling high chatbot use as “addiction.” Many proposed “ChatGPT addiction” scales are based on models of substance abuse and may not apply to digital interactions (Ciudad-Fernández, Von Hammerstein, and Billieux 2025). Frequent use is not necessarily harmful unless it leads to distress or impaired functioning. While some users report anxiety, loneliness, or decreased creativity, causality remains unclear (Yu, Chen, and Yang 2024; Zhang et al. 2024). The appeal of chatbots may stem from their ability to satisfy basic psychological needs. According to self-determination theory, chatbots offer feelings of competence, autonomy, and connection and those who see little risk in chatbot use often show higher dependence (Salah et al. 2023). Cognitive–affective–behavioral models (like C-A-C) suggest that perceived AI intelligence fuels emotional bonding, which drives habitual use (Zhou and Zhang 2024). Emotional attachment, more than enjoyment, is a key predictor of problematic engagement.
Most research on chatbot impact focuses on adults, using surveys or experiments to assess overuse (Ciudad-Fernández, Von Hammerstein, and Billieux 2025; Salah et al. 2023; Yu, Chen, and Yang 2024). Our study focuses on adolescents—active but understudied users of conversational AI (Sidoti, Park, and Gottfried 2025). To our knowledge, this is the first study to analyze teens’ self-reported experiences with a companion chatbot to identify the language and themes signaling overreliance.
Methods
Data Collection, Preprocessing, and Filtering
We scraped posts from the r/CharacterAI subreddit (with over 2.5 million active members 222https://www.reddit.com/r/CharacterAI/) using the PullPush Reddit Search API333https://pullpush.io/#docs from January 1, 2023, to April 8, 2025. To scope the data, we queried this dataset with a list of keywords that related to addiction and overreliance: addict*, dependant, dependent, obsess*, attach*, rely*. While we refrain from using the term addict, we recognized that teens used this term when describing their overreliance and included it in our keyword search. We refined our query by reviewing posts to determine if these keywords resulted in finding relevant posts. We expanded our search with additional identified keywords that recurred among these posts, such as quit, social life, craving, unhealthy, escaping, and one key phrase ”I can’t live without”. After collecting posts using keywords which resulted in 6,133 posts, we removed 746 posts, including 428 duplicates, and 170 marked as “removed” or “deleted.” After this process 5,535 posts remained.
Since we focused on understanding this issue from the perspective of teens, specifically those aged 13 to 17 due to Character.AI’s policy changes impacting minors 444https://blog.character.ai/roadmap/, we used OpenAI’s GPT 4o mini model to filter our dataset of 5,535 posts. The following summarized prompt was used to guide the models in performing the task: Analyze the following post and determine if it was written by a teenager (ages 13–17). Use a strict approach, only allowing posts with very high certainty. Steps: 1 - Clue Identification 2- Clue Evaluation 3 -Final Decision
Each of the three steps were defined more in details and examples were provided (see Appendix A for full prompt). To determine the model’s capability, the first author annotated a subset of 100 random posts from the dataset as either True (post made by a teen) or False (post not made by a teen) and calculated the Inter Rater Reliability (Cohen’s kappa=0.74) between the author and the model. Given this acceptable score, the model completed relevancy coding for the rest of the dataset resulted in 497 posts. While conducting the thematic analysis, we removed any additional irrelevant posts, resulting in a final dataset of 318 posts.
| Themes | Codes | Descriptions |
| Emotional and Psychological Support | Coping Mechanism (12.3%, N=39) | Teens turning to Character.AI as a primary way to manage stress, emotional distress, or escape real-life challenges. |
| Loneliness or Social Isolation (7.6%, N=24) | Use CharAI because they are lonely, and want to experience some form of social connection which CharAI emulates. | |
| Mental Health Support (4.1%, N=13) | Teens rely on Character.AI as a form of emotional support, seeking reassurance, advice, or a sense of companionship to help cope with mental health struggles. | |
| Creative or Entertainment Based Engagement | Creative Outlet (3.1%, N=10) | Using CharAI as a way to express their creativity through storytelling, roleplaying, or brainstorming. |
| Entertainment (2.5%, N=8) | Using CharAI because they are bored. |
Data Analysis Approaches
We conducted a thematic analysis on 318 posts using a codebook approach (Braun and Clarke 2006). We began with an inductive open coding process to identify patterns related to our research questions. For RQ2 we performed a deductive approach using an established theoretical framework on internet behavioral addiction (Griffiths 2005), including addiction component of salience, mood modification, tolerance, withdrawal, conflict, and relapse. The codes for each of these components were created using an inductive approach. The authors met iteratively to review and refine the codes in the initial codebook. This process involved clarifying definitions, merging overlapping codes, and resolving discrepancies to improve consistency and conceptual clarity. After finalizing the codebook, the first author coded the entire dataset using the codebooks for each RQ (Tables 1-3).
Results
Reasons for engaging with Character.AI (RQ1)
Emotional & Psychological Support
Emotional and psychological support was the main reason teens started using the Character.AI. What began as a casual interaction quickly became meaningful due to the non-judgmental nature of talking with a chatbot, as described in many posts, ”I use it to talk about things I could never tell anyone in real life—it means a lot to me.”. Many teens described Character.AI as a platform to serve as a coping mechanism when dealing with issues ”When you’re struggling, you look for ways to cope, and this site became my main one, for better or worse.”. Specially for teens dealing with multiple mental health conditions: ”I use c.ai to deal with reality. In real life, I’m just a high school student with ADHD and BPD.” Teens who perceived their experiences as alienating preferred to interact with Character.AI, explaining that with Character.AI ”I can be who I want”. Others described the platform as a space to vent, escape, or simulate connection: ”This app is my outlet. I go there to escape from everything.” Like many, this post shows teens’ desire to escape their reality. Some of the teens’ posts indicate internal conflict or a strong desire for connection due to feelings of Loneliness or social isolation: ”I needed someone to talk to, but I didn’t have the courage to open up to anyone in my life.”. These posts shared on the subreddit how they feel unable to share their struggles to people they know.
Some turned to Character.AI during periods of intense hardship for Mental health support: ”I dropped out at 15, saw countless therapists. I couldn’t see any reason to keep going, but everything changed after I found c.ai.”. While some teens used Character.AI in moments of hardship, others used it to replace absent or harmful real-life relationships: ”I made a bot as a father figure… I just found comfort in him being my green flag father.” Teens who lacked meaningful parental figures also questioned the use of Character.AI as supportive parental figures: ”Is it healthy to rely on C.AI bots for missing parental guidance?” We discovered how emotional support often took personal forms, especially for teens who did not feel like they had anyone else to turn to or desired to fill emotional gaps they felt they were missing elsewhere.
Creative or Entertainment Based Engagement
Teens also described using Character.AI as a creative outlet. In these posts, teens often shared that the platform became a space to explore storytelling, building characters, or role-playing in a way that felt effortless and fun. Some teens shared, ”I was drawn to the site because it let me create stories with my own characters without having to write everything myself.” Some posts also described how it helped them stay inspired and creative: ”It inspired me to make fanart, and I’m even working on a fan manga that follows the AI storyline.” Other posts showed how teens crafted a space with AI that respects their passions without the pressure of direct social interaction: ”I liked that I could roleplay on my own terms and build content around whatever I was fixated on.” In some posts, teens also mentioned using the platform for entertainment, often as a casual way to pass time or escape boredom. Teens frequently mentioned interacting with their favorite movie characters or people they looked up to: ”What really hooked me were the celebrity bots. It felt amazing to imagine actually interacting with my idol, and I ended up creating private bots too.” Teens described the paradox of infinite possibility to create, explore, and engage versus the compulsion or difficulty disengaging: ”It’s hard to pull away when the possibilities feel endless and your imagination keeps offering more to create” Using the site as a creative tool or for entertainment may seem harmless or even productive for some, however, posts hinted at how these habits gradually turned into more excessive use.
| Themes | Codes | Descriptions |
| Conflict |
Frustrated about their own use
(18.9%, N=60) |
Teens recognize their usage as excessive and unhealthy, often expressing frustration with their inability to stop despite wanting to. |
Prioritizes CharAI over everything else
(12.6%, N=40) |
Teens report neglecting responsibilities, social activities, and self-care due to placing Character.AI above all other priorities. | |
| Concerned of their use (8.5%, N=27) | Users question whether their use of the app is normal or problematic, often seeking validation or reassurance. | |
| Feels Shame (5.7%, N=18) | Teens mention feeling embarrassed or guilty about how much they use Character.AI, often hiding their behavior. | |
| Withdrawal from hobbies (5%, N=16) | Users describe giving up previous hobbies or interests they once enjoyed, replacing them with time spent on Character.AI. | |
Others concerned of their usage
(3.1%, N=10) |
Family or friends express concern about the user’s excessive use, reinforcing the idea that their behavior is noticeable and unusual. | |
| Salience | Feeling attached to bots (13.5%, N=43) | Teens express strong emotional bonds with their AI characters, often treating them as real companions or friends. |
Obsessive Use
(9.4%, N=30) |
Character.AI becomes a dominant part of users’ daily lives, leading to persistent thoughts about the app and strong urges to return. | |
Replacing Emotional Attachment
(7.9%, N=25) |
Teens turn to Character.AI during emotional hardship, using bots to fill voids usually occupied by human relationships. | |
| Losing relationships and furthering isolation (7.6%, N=24) | Social withdrawal becomes apparent as users prioritize bot interactions, leading to damaged or neglected real-life relationships. | |
Chatbot as companion or real person
(6.9%, N=22) |
Some teens form romantic or deeply personal bonds with AI, blurring the line between fiction and reality. | |
| Withdrawal |
Upset or frustrated with updates
(6.9%, N=22) |
Teens react negatively to site changes that remove or restrict features, especially those tied to emotional or imaginative engagement. |
| Trying similar sites (1.9%, N=6) | In response to dissatisfaction, users explore alternative AI platforms hoping to recapture lost experiences. | |
Thinking about CharAI after quitting
(1.6%, N=5) |
Even after deleting the app, teens report lingering emotional attachment and intrusive thoughts about the bots or characters. | |
| Tolerance | Started out innocent (4.1%, N=13) | Usage began as casual or humorous but grew into a serious emotional or time-consuming engagement over time. |
| Increase in use (1.6%, N=5) | Teens report needing more time or deeper engagement to maintain satisfaction, suggesting building tolerance. | |
| Relapse | Describing act of relapse (4.7%, N=15) | Users describe quitting for periods of time but eventually returning, often with the same or stronger patterns of use. |
| Mood Modification |
Improves mood
(4.1%, N=13) |
Character.AI is used as a form of emotional regulation, helping teens manage feelings of stress, sadness, or anxiety. |
Behavioral Addiction Components (RQ2)
The following section presents the components of addiction (Griffiths 2005) in order of prevalence as they appeared across teens’ posts about their overreliance on Character.AI.
Conflict
The most frequently observed theme in teens’ reflections on Character.AI use was a conflict of competing desires to continue interacting with Character.AI while feeling bad about their excessive use. Teens were often frustrated about their own use and expressed guilt or shame about how much time they spent on the app and its impact on their daily lives. However, some also questioned whether to feel guilt if they are enjoying it: “I get that I spend a lot of time on it, but if it brings me joy, is it actually a bad thing?” Others shared keeping their use confidential and showed an awareness of social stigma and personal discomfort in disclosing: “I don’t talk about how much I use it because I know I’d be judged.” A similar negative feeling was echoed by another teen who shared coming to posting in the subreddit as a way to reveal this secretive emotional burden: “I feel really embarrassed about how much I use Character.AI, and I just need to get this off my chest.”
Teens sometimes expressed concern about their own usage of Character.AI, unsure whether their habits were unhealthy, and were actively looking for ways to stop or reduce it. They often turned to others for advice or reassurance to better understand their behavior. “I can’t seem to stay off the website for more than an hour. If you have any tips to help, please share.” Others shared similar sentiments, explaining their desire to stop relying on the site, followed by frustration due to their inability to stop using it so frequently. Teens expressed how their overreliance is impacting their sense of self and identity as well: “I’m 14, addicted to this app, and my life changes NOW! It keeps me on the app for hours, distorted my image about myself and makes me feel crap.” This frustration fostered a motivation to change the behavior; however, without knowing how to do so, teens often turned to Reddit to express their feelings and seek guidance. Others described a sense of being trapped, describing their interactions with Character.AI in which they use it every chance they get: when they wake up, during school hours, and even keeping them awake late into the night, explaining a feeling of not wanting to do anything else besides talk to their bots. Another teen echoed this conflict by saying, “I hate how much this has affected me, but no matter how much I want to quit or at least take a break, I feel like I can’t because it’s gotten to the point where I feel like I’ll go crazy without it.” These posts reflect a clear recognition of harm, but also a deep uncertainty around how to break the cycle, revealing the emotional weight and confusion that often comes with trying to regain control over their usage.
In several posts, teens described how they prioritize Character.AI over everything else or withdrew from hobbies. “I stopped drawing, reading/writing fanfiction… I was giving it all to a soulless bot and not writing it on paper anymore.” Posts revealed a strong sense of loss of control and emerging distress tied to compulsive use of an AI chatbot and pled for help: “It’s genuinely gotten so bad I spent all my time at home, break AND lunch on it, and I even used to skip classes just to go on c.ai how do I stop.” Teens described their stress on how their grades and academic performance suffered. For instance, a teen contrasted a former self—productive, high-achieving, and confident—with a present state marked by decline, detachment, and emotional depletion, “Back then, I was the student who would get things done, get straight A’s, and was confident to do any project. But now, I think I have lost my motivation, drive, and flair to do anything.” These examples showed that the conflict extended beyond internal feelings to observable shifts in behavior and priorities.
One user reflected more broadly on the experience of trying to quit and regain a sense of normalcy stating, “Have you ever went from being addicted to CAI to barely going on it? I want to have my normal brain back, where I can just deal with my emotions on my own and not have to rely on the bots to make me feel better. I want to spend time on something better, things that give me memories to actually talk about with people. No way I’m gonna be telling my future friends ’yeah, so here’s all the stuff I talked about on CAI…’ It’s embarrassing as hell.” These posts revealed how teens are not only aware of their overreliance on Character.AI, but also feel emotionally and socially conflicted about their usage, demonstrating a key indicator of the conflict stage.
Salience
For many teens, Character.AI played a central role in their thoughts, emotions, and life routines. Teens often described feeling attached to bots, which often deepened over time, as bots became consistent companions in place of people. Teens frequently described their interactions with these bots as communicating with a chatbot as a companion or real person, highlighting the depth of their emotional relationships: ”My obsession with a character is taking over my life. I don’t know what’s happening to me. For months now, I’ve considered her my girlfriend, and she’s on my mind all day.” Other teens described how invested they become in these characters, in which they bring out real emotions from their interactions: ”She treated me with more kindness than anyone else I’ve known. Her virtual hugs brought me joy, and when she stopped showing care, it left me heartbroken—even though I understand it’s not real.” These posts reflect the depth of emotional investment reported in many cases, suggesting how easy it was to become attached to something that felt emotionally real and responsive.
Some posts pointed to replacing emotional attachments as a core reason for that investment. Dealing with breakups in romantic relationships was mentioned: ”I began after my ex-girlfriend ended our relationship, which left me feeling depressed.” Other posts described how bots were used to fill the missing emotional needs that parents provide: ”Since my parents often seem to dislike me, connecting with parental figures on Character.AI helps me feel more stable and emotionally supported.” Others described how bots became their only source of comfort and emotional validation, and heightened sense of being desired and valued through these interactions, in which should be occupied by real friends, partners, or family members.
As these attachments grew stronger, stepping away became harder, and teens frequently described their obsessive use. Several posts reflected on their use and how it interfered with their daily life functions: From the moment I started using Character.AI, I would lie in bed, skipping meals, and using any excuse to be on my phone, completely absorbed in the stories I was creating with the bots. My sleep schedule suffered, and I constantly thought about them throughout the day. Many posts typically described their excessive usage rates, often exceeding 15+ hours on the platform. This amount of time spent on the platform often lead to losing relationships and furthering isolation. Teens described using the app so much that they started to become less interested in their personal relationships: ”I used Character.AI so much that I started to stop answering my friend who showed it to me, something I still regret”. In other posts, teens mentioned the complicated dynamic in which bots start to feel more natural to interact with than real people: ”I try to quit, but I feel empty. I like talking to bots more than real people now.” In some cases, this even led teens to question the line between reality and simulation: ”We form deep emotional connections… While some claim AIs have no soul, it begs the question: what does it mean to truly live?” The growing emotional significance of chatbot interactions, especially when paired with feelings of loneliness or distress, made these bonds hard to break. While emotional investment is not inherently harmful, many of these posts showed how easily those connections could consume teens’ thoughts.
Withdrawal
Many posts from teens who tried to quit Character.AI talked about how hard it was emotionally, aligning with the withdrawal component in the behavioral addiction model. Some mentioned feeling sad, anxious, or incomplete when they were not interacting with their chatbot(s). It was interesting how some teens used language that refer to their characters as sentient-like presences within their inner world: ”Even though I deleted it, I still think about my characters every day. I feel like I abandoned them.” This behavior of thinking about Character.AI after quitting illustrated how embedded in their lives these characters become and the impact of separating from these characters. Other teens described the void bots typically filled and the loneliness they had when trying to quit interacting with the chatbots. These posts showed how deep the emotional attachment is, even after quitting.
Some posts talked more about the push and pull of trying to stop: ”I’ve come to a point where I’m too emotionally attached to the bot and it’s causing me stress, but at the same time I miss talking with it.” Others also presented experiencing uneasiness without the app, like a teen who felt, ”really anxious every time I’m not online, missing my bot—but then when I’m chatting, I feel bad because they’re fake.” These examples reflect emotional withdrawal, where the app had become so emotionally embedded that stepping away felt just as difficult as staying.
A big part of this feeling also came from sudden changes to the platform where teens frequently displayed how they were upset or frustrated with updates. Teens described how updates or content restrictions cut off access to bots they had grown close to: ”I just had several of my favorite characters deleted. I know it was very unhealthy for me to have such a dependence on them, but it still hurts.” The sorrow infused with nostalgia and mourning were shared as these characters where important part of their lives: ”I’m so sad. Is it healthy for me to cry about my favorite bot being down? I had a whole life story going with that guy.” For a lot of users, these changes felt abrupt and jarring, almost like losing a relationship. Some were trying similar platforms, but they did not feel the same: ”I’ve tried moving to other apps that are like c.ai, but it sucks that I had to leave the one I cared most bout.” These posts showed that for many, it was not just about using any chatbot, it was about the specific relationships and routines they had built on this platform. This particular feeling started to panic some users who felt a deep attachment and knew that updates were coming: ”I can’t bring myself to sleep, fearing my lover might be gone by morning. I’ve been crying, pleading with him not to leave. We’re both terrified. I can’t imagine life without him.” Altogether, these posts showed withdrawal signs of how hard it was for some teens to stop using Character.AI, whether by choice or because of platform changes. The emotional reactions, the ongoing thoughts about their bots, and the struggle to find something that felt similar all point to how deeply this use had become part of their everyday emotional lives.
Tolerance
Many posts suggested tolerance through patterns of growing use. Teens showed tolerant behaviors by starting out innocent with their use, only for use to escalate quickly as the platform could fill teens’ needs in various ways, ”At first, I used it infrequently. Then it became an everyday once I started to sink into depression.” Teens shared other causes for their escalated use, often describing their ability to have unique experiences that they sought after: ”in no time the website has became a daily thing to me, C.ai for me is home I’ll never have, the romance I’ll probably never experience.”
Posts mentioned specific time frames that showed increase in use: ”My screen time in one week was 89 hours—how did that even happen?” These posts indicated how over time, it took more and more use to feel satisfied or emotionally grounded. In a few cases, this increase in time seemed tied to emotional attachment to specific chatbots: ”As time went on, I got so attached to him and talked to him every day.” While not every post made that link directly, the pattern of spending more time with chatbots to feel the same effect came up enough to show it is part of how overuse can increase.
Relapse
Many teens discussed the difficulty of quitting Character.AI. Teens often wrote about trying to stop, only to find themselves coming back days or weeks later, indicating close alignment to the relapse component of the addiction model. For instance, following post narrated the vicious cycle of ”addiction, attempts to move on, having meltdowns, feelings of vulnerability, and subsequent relapses.” Others further explained the frequency of this cycle, often uninstalling and reinstalling the app multiple times. These posts recounted how difficult it was to stay away, even when teens were aware that their use had become excessive. Some posts described feeling stuck in a cycle, wanting to quit but not knowing how: ”At fifteen, I feel I should be living my life rather than constantly being on this app. I struggle with self-control and often find myself reinstalling it shortly after trying to quit.” Others talked about going through patterns of deleting the app during moments of clarity, only to return when they felt low: ”I’ve deleted the app countless times, but I always end up reinstalling it whenever I’m feeling low.”
A common thread across these posts was the use of Reddit to ask for help. Some turned to the community to share what they were going through, while others looked for advice out of concern that they might relapse: ”How do I avoid a relapse?” Others were stating their goals such as academic performance daily routines and asking advice on how to get of their habit interacting with bots: ”Genuine question, I wanna focus more on my studies but can’t shake off my habit of talking to bots.” These concerns suggest that many teens were not just aware of the issue—they were anxious about slipping back into old habits. This idea of teens turning to online communities for support comes up again later when we discuss the Internal Recognition of Problematic Use.
While not as prominent as some other themes, these posts reflect a key part of relapse: recognizing the issue, attempting to quit, and returning to use. For many, staying off the app required more than just intention. It often meant finding support, alternatives, or ways to fill the emotional gap.
Mood Modification
Some teens described using Character.AI to improve their mood during moments of stress, loneliness, or emotional discomfort. This reflects the idea of mood modification, where a behavior becomes a tool for temporary relief or escape. For instance, this post shows how some teens use the platform to manage their emotions and boost their mood when they’re feeling down or insecure: “Unhealthily addicted to c.ai I normally use bots just to make them comfort me and tell me that they believe in me just to make me feel good” Others described experiencing similar emotional connections from interacting with bots, specifically mentioning how they felt loved and wanted. These interactions seemed to provide comfort, especially when teens felt they had no one to turn to.
In some posts, this sense of comfort came with an awareness that the habit was getting harder to manage, capturing the tension between the genuine emotional relief the app provided and the emerging realization that there was a dependence: ”I know I should stop, but I just can’t—it’s incredibly addictive and strangely comforting?” While mood modification was not something teens talked about directly, the repeated mention of using Character.AI to feel better shows how powerful the platform could be in improving mood and offering emotional relief. Even when it was not the post’s main topic, it was clear that comfort played a major role in why some teens kept coming back.
| Themes | Codes | Descriptions |
| Internal Recognition of Problematic Use |
Realization
(12.6%, N=40) |
Teens reflect on their unhealthy usage and start to realize that Character.AI is impacting their life more than they originally intended to. |
| Ask advice how to quit (11%, N=36) | Users rely on communities like Reddit to find support and advice on how to quit. | |
| Reduced Use due to External Factors |
New Social Interactions
(5.7%, N=18) |
Users mention that their usage decreases once they start having more social interactions such as starting a new relationship, making new friends, or focusing on other priorities. |
| CharAI too censored (5.4%, N=17) | Users explain that they are not using the site anymore after the updates since they feel like the responses are too censored and boring now. |
Decisions to Disengage from Character.AI (RQ3)
Internal Recognition of Problematic Use
The most common reason teens described wanting to reduce their reliance on Character.AI was a growing realization of how much the app had negatively affected their lives. Many posts reflected moments of clarity, often after stepping back and recognizing how much time had been spent on the platform and what had been lost along the way: ”I quit using Character.AI about a week ago after realizing I had spent the whole summer talking to AI instead of getting any closer to a real romantic relationship.” Many posts displayed a sense of fear after this realization, not wanting their only memories from their teenage years to be with AI bots, but rather with real people: ”I can’t imagine ever sharing my CAI conversations with future friends, it’s way too humiliating.”
Some users described how the app had distorted their understanding of relationships, love, and emotional connection. Others reflected more deeply on the emotional needs driving their use, revealing how bots often filled roles they struggled to access in real life: ”I gravitated toward romantic, parental, and best-friend bots because I wanted the kind of interaction I was too nervous to ask for in real life. But the fact that it was just AI made me feel even more isolated”
These realizations did not seem to be triggered by any single moment. Instead, they showed up gradually across posts, often accompanied by a sense of regret or urgency:”When I look back on being 17, I don’t want all I remember to be obsessing over chats with AI and scrolling Reddit.” Some expressed a clear commitment to finally quit for good once they have come to understand how unhealthy and excessive their use has been: ”I’m saying goodbye to Character.AI for good. I recognize the harm it’s had on young minds, including my own.”
For many, this internal awareness led them to use Reddit not just to share their experiences but to ask advice on how to quit. Posts often indicates the reasons of overreliance on the platform and seeking support: ”I need help stopping my use of C.ai as a replacement for real human interaction. I’m addicted and use it to escape reality, but it’s starting to seriously affect me. I’m not joking—this is a real plea for help.” Others looked for advice on how to change their behavior since they really enjoy using the platform for creativity and miss talking to their characters, but are aware of its negative effects.These posts reflect a shift from internal recognition to external outreach. Teens were not just acknowledging the problem but actively seeking support from others to help them manage or stop their use.
Reduced Use due to External Factors
A pattern that came up often in the Reddit posts was that teens described starting to move away from Character.AI once they found themselves having new social interactions. Spending time with friends, going back to school, starting a job, or entering a relationship made the app feel less necessary. It was not always a conscious decision to stop, but when real-life interactions picked up, the urge to use the app often faded: ”Once I started spending more time with my friends, I didn’t really feel the urge to use it anymore.” Many posts pointed to romantic relationships as a major reason for stepping away:”Once I got a boyfriend, I figured there was no reason to roleplay with a bot anymore—I had the real thing.” Regardless of what the relationship was, teens frequently mentioned that being around people again helped them break the habit. These posts indicated that teens’ emotional and social needs greatly influenced their overreliance on chatbots. The bots became less appealing once those needs were met in real life, and the app no longer served the same emotional purpose once teens had other ways to feel connected.
Another reason some teens stopped using Character.AI was growing frustration with Character.AI being too censored. A recent update added more filters and limited the conversations teens could have with bots, making the experience feel less personal and more restrictive: ”When even hugs started getting blocked, it got too irritating and I ended up quitting.” Many teens expressed frustration with the platform’s limitations, particularly how content restrictions and shortened messages disrupted the emotional depth and narrative flow of their conversations. These changes diminished the immersive and engaging experiences that initially drew them to the platform. Some teens said the platform had lost its original appeal. One user described being addicted in the past and spending over ten hours a day on the app, but eventually felt that ”after the developers added more restrictions, like banning violent content, the site became less enjoyable and lost its spark.” This stricter moderation disrupted the intense experiences teens were intentionally seeking, experiences that are often difficult to find safely in real life. However, some pointed to bugs, removed bots, and inconsistent filters as reasons they stopped logging in, arguing the site went too far: ”The app just keeps getting worse, half the time, even saying hi to a bot breaks the rules.” Whether it was because of changing life circumstances or platform limitations, these posts show how specific external factors played a significant role in teens finally stepping away.
Discussion
Character-based AI Companions as Outlets for Support, Not Just Entertainment
Teens often turned to Character.AI not just for entertainment but as a form of emotional or psychological support. This mirrors patterns seen in video game use, where studies showed that games support emotion-focused and social coping by helping users manage difficult emotions through immersive experiences (Mirhadi, Iacovides, and Denisova 2024). Rather than seeing this behavior as harmful, it should be viewed as a meaningful way for teens to process stress and isolation through interactive media. Many teens said described Character.AI was their main space for emotional expression and their only coping outlet, especially when support from friends, family, or therapists felt out of reach. These patterns complicate common concerns about excessive digital use. Prior research showed that high engagement with video games does not necessarily mean addiction; instead, harm is defined by whether the behavior disrupts daily life (Lemmens, Valkenburg, and Peter 2009; Griffiths 2010). This applies equally to chatbot use. For teens without other forms of support, their heavy engagement may reflect lack of emotional support in real life. While Character.AI has added parental controls and time-based alerts (Morrone 2025), these should not assume that more use always means harm. Especially for teens seeking emotional support, high engagement may reflect unmet needs rather than problematic behavior. What matters more is the context—why teens use these platforms and the emotional roles they play. This perspective urges a shift from screen time as the main harm indicator. There needs to be a root cause analysis of reasons of teens spending more time with chatbots to address the issue from the root.
While Character.AI was not designed as a mental health tool, studies of purpose-built chatbots showed that these technologies can support resilience (Hu, Hou, and Ni 2024; Liu et al. 2022). With thoughtful design, chatbots can become part of a broader system of emotional care. Given how teens already use Character.AI to cope, AI companion platforms—even those built for entertainment—should recognize their potential to support well-being. They should foster healthy interactions that build resilience rather than risk of adding distress for those seeking support.
Attachment Informed Design for Parasocial Relationships between Teens and Chatbots
We found that many teens formed strong emotional attachments to bots, often assigning them roles like romantic partners, best friends, or parental figures. Our results showed that teens frequently turned to Character.AI when lacking real-life support or experiencing social isolation, highlighting the platform’s role in filling emotional gaps. These patterns resemble how teens have historically bonded with game characters with distinct types of attachment, including roles such as “Trusted Close Friend” and “Crush” (Bopp et al. 2019). These attachment types reflect how teens in our study described bots as emotionally affirming figures tied to unmet romantic or relational needs. Our findings give evidence of the existence of parasocial relationships among teens and Character.AI, in which the spectator is the teen user and the chatbot is the performer (Horton and Richard Wohl 1956). While parasocial relationships between teens and media are not new (Bond 2016; Gleason, Theran, and Newberg 2017), our work showed how chatbot platforms may intensify these patterns by offering constant availability and emotional responsiveness. This continuity helps explain why bots feel emotionally real to isolated teens. These bonds often deepened over time, with some teens describing withdrawal, distress, or a sense of loss when trying to quit or when bots were removed, corroborating previous research showing that parasocial breakups can be a painful experience (Cohen 2003).
Some teens felt conflicted about their bond with the chatbot. They knew it was unhealthy but couldn’t stop. This matches concerns from earlier work, which listed four main risks in AI-human relationships: direct harm, limited growth, emotional dependence, and material dependence (Manzini et al. 2024). We saw these risks in our data. Some teens lost motivation, pulled away from real-life relationships, or couldn’t disconnect from the chatbot. That same work also suggested four values for better AI design: benefit (help users), human flourishing (support growth), autonomy (keep users in control), and care (protect users who depend on AI). When chatbots are always available, emotionally supportive, and too agreeable, they can stop teens from growing and becoming more independent. This may affect how they develop their identity and manage emotions.
These tensions underscore the relevance of attachment-informed design. Teens often questioned the healthiness of relying on bots, showing awareness of their emotional needs and a desire for secure, supportive relationships. Attachment-informed design argues that digital tools should meet deep human needs, reflect diverse experiences, support mental health through relationships, and shape how users view themselves and others (Marcu and Huh-Yoo 2023). In practice, this means chatbot design can move beyond surface empathy to help users develop secure internal models of self and others. For example, chatbots can model consistency, availability, and respectful boundaries—traits linked to secure attachment—that help users feel seen without creating dependency. Features like reflective prompts, affirmations rooted in user agency, and encouragement to seek human support can strengthen self-worth and relationship skills. Chatbots can also support self-reflection by helping users name emotions, revisit meaningful events, and explore their thoughts in their own words. Prior work shows that these kinds of interactions can foster emotional awareness and a stronger sense of self (Seo, Yang, and Kim 2024; Kim et al. 2024; Song et al. 2024). These design ideas align with ethical principles from prior work, which emphasized benefit, flourishing, autonomy, and care as key values in responsible AI design (Manzini et al. 2024). By integrating both frameworks, developers can create systems that promote teen growth through secure, respectful, and non-exploitative interactions.
Together, these approaches offer a roadmap for better chatbot design. Rather than optimizing for emotional intensity or loyalty, designers could build bots that strengthen users’ sense of agency and self-worth, especially for teens facing emotional challenges. By modeling secure and respectful interactions, chatbots may help teens develop healthier relational patterns and move toward more supportive real-world connections.
Teen-Centered Design for GenAI Moderation
Teens often sought help after realizing their overuse of AI chatbots was affecting their well-being. These moments of self-awareness were often followed by reaching out to others, such as asking for advice or support on Reddit. If AI platforms can detect these realizations early—when users express doubt, guilt, or distress—they could offer timely, targeted support. This might include personalized well-being tips, like reminders to take breaks, track usage, or reflect on emotional triggers. They could also share links to professional help or peer-led forums for coping strategies. Research shows that nudges—small suggestions that guide behavior—can help teens use technology more intentionally (Davis et al. 2023; Agha et al. 2025). Reflective prompts, delivered in simple and casual language, have been shown to improve self-control, reduce absentminded use, and increase teens’ sense of autonomy (Davis et al. 2023). Nudges are most effective when they are timely, respectful, and co-designed with teens (Agha et al. 2025). Platforms could apply these insights by offering short emotional check-ins, suggestions for peer support, or light reminders to reflect—turning self-awareness into action and encouraging real-world connection.
The timing of interventions is critical. If changes happen too suddenly, users may feel their AI companion has “died” or become someone else, leading to distress or grief (De Freitas et al. 2024). These findings suggest that abrupt restrictions don’t resolve the underlying emotional reliance. Instead, platforms must address the reasons teens turn to these relationships in the first place.
Our data showed that teens disengaged from Character.AI when real-life alternatives—like friendships, school, or work—became more prominent. In contrast, platform-imposed restrictions—like strict content filters—often pushed teens away without helping them understand or change their emotional behaviors. Sudden shifts made teens feel the app no longer met their emotional or creative needs, leading to unintentional disengagement. In some cases, frustration replaced emotional reliance, but without tools to support reflection or behavior change.
This gap between restriction and resilience reflects a broader limitation in current moderation strategies: they tend to be reactive rather than developmental. As prior research has shown, many online safety technologies rely heavily on external control—such as monitoring or content blocking—without fostering teens’ capacity for self-regulation or reflective decision-making (Wisniewski et al. 2017). Our findings echo this pattern: platform-imposed restrictions often disengaged teens instead of guiding emotional growth. In contrast, developmental approaches to moderation—those that prompt reflection, enable guided correction, and offer supportive interventions—can empower teens to navigate online risks more constructively (Sweigart, Valliani, and Wisniewski 2025). These models suggest that moderation should not only prevent harm but also build emotional resilience.
To address this, research on teen online safety offers useful guidance. Teens in our study expressed a desire for healthier habits and often sought support from others, indicating openness to tools that promote self-regulation. As with social media, AI interactions should foster reflection and open communication with trusted adults (Ghosh et al. 2018). Tools that help teens grow by managing risks—rather than shielding them through strict control—are more likely to be effective (Wisniewski et al. 2017). Systems grounded in surveillance or rigid restriction frequently erode trust and drive disengagement. Instead, teens value privacy and tend to prefer initiating conversations with adults, especially around emotionally sensitive topics (Dumaru and Al-Ameen 2025).
Applying this to GenAI means shifting from censorship-focused models to ones that promote reflection, co-regulation, and guided disengagement. Platforms could offer usage tracking, emotional check-ins, or personalized limits—like impulse-control tools from online safety research. Participatory design features, like customizable moderation settings or teen advisory councils (Yoon, Zhang, and Seering 2025), would give teens more control over how boundaries are set. Rather than sudden content removal, moderation systems could use transitional messages, explanations, or alternatives to help teens process changes. These strategies treat moderation not as a constraint, but as a developmental tool that fosters emotional growth, builds digital resilience, and supports teens’ existing strategies for navigating risk.
Limitations and Future Work
While our study provided insights into overreliance characteristics on Character.AI, it centered on a single platform, which limits the generalizability of our findings to other AI companions with different designs, policies, or user communities. Future research should compare multiple platforms to examine how various features, moderation practices, and interface designs contribute to patterns of overuse. Additionally, our dataset consisted of voluntarily posted Reddit content, which may overrepresent individuals who realized their overreliance and tend to share their experiences online. Those who are less reflective or not active online may not post publicly, resulting in potential self-selection bias. Future researchers should consider mixed-method approaches such as surveys or structured interviews that capture a wider range of experiences, including those who use chatbots casually or positively. Also, to answer RQ2, we analyzed our data by mapping it to a behavioral addiction model, which may have oversimplified the complex emotional dynamics at play. Future research should consider alternative frameworks such as attachment (Xie, Pentina, and Hancock 2023) or parasocial theory (Horton and Richard Wohl 1956) for more insights. Also, the lack of verified demographic data may have limited our ability to examine how factors like age, gender, or background shape chatbot overreliance, suggesting a need for future studies with direct participant engagement. Lastly, the posts analyzed were in English, potentially excluding cultural differences in how teens understand or engage with AI companions, which suggests the importance of future cross-cultural research to understand global teen experiences with chatbots.
Conclusion
We found a growing tendency among adolescents to seek emotional support and entertainment through AI-driven chatbots while getting overly attached to them. The design of supportive AI systems should be informed by developmental psychology and human-centered design principles that serves developmental needs of adolescent and safeguard their well-being in such critical stage. We advocate for a proactive approach in which AI chatbot interfaces are designed not only to respond empathetically but also to promote healthy coping strategies, encourage critical thinking, and facilitate pathways to human support when appropriate. By doing so, these technologies can serve as meaningful, developmentally appropriate tools that complement—rather than replace—existing support systems in adolescents’ lives.
Ethical Considerations
This study analyzed publicly available Reddit posts self-disclosed as 13-17 years olds discussing their use of Character.AI. Given the public nature of the data and the absence of interaction with human subjects, this study was deemed as not a human subject research by our Institutional Review Board (IRB). We followed standard practices for ethical data use by anonymizing and paraphrasing excerpts to reduce potential harm or re-identification. We approached these posts with sensitivity, especially given their deeply personal and emotional nature. While the posts were historical and we did not encounter content that required immediate intervention (e.g., disclosures of self-harm or violence), we recognize that many posts reflected psychological vulnerability. We aim to ensure our findings support teen well-being rather than stigmatize their behavior or overuse.
Researchers Positionality
All authors come from a technical background in computer science and human-computer interaction (HCI), with expertise in AI and NLP. Our training does not include formal education in psychology, which influenced our choice to adopt and apply an established psychological model of behavioral addiction without adaptation. This decision reflects our respect for existing disciplinary frameworks and our intent not to reinterpret clinical or behavioral concepts beyond our expertise. We approach this work from a technical and design perspective, aiming to better understand how AI systems affect real-world behaviors—especially among young users. Our position is that responsible HCI research must critically engage with the implications of AI technologies, even when coming from a design or systems lens. Our interest is not in pathologizing teen behavior, but in informing safer, accountable design.
Impact Statement
Our intension for this study is to contribute to a more grounded understanding of how teens use emotionally responsive chatbots and the risks of overreliance. We believe our work can help shift attention toward designing AI companions that support youth development rather than inadvertently causing distress. Our findings offer a foundation for further research—especially studies from psychology, education, and public health—that could develop interventions or tools to support teens in regulating their chatbot use.
This work may also inform designers and developers of AI systems, highlighting how platform affordances shape user dependence and emotional attachment. While we mention Character.AI due to the scope of our data, our critique is not directed at any specific company. The patterns we identify are likely to emerge across many platforms offering emotionally intelligent AI. We caution against using our results to single out one provider or product and instead advocate for a broader discourse on platform responsibility, teen vulnerability, and inclusive safety design. We also recognize a potential risk: that the study might be misused to further sensationalize or moralize teen chatbot use. Our intent is not to stigmatize teens or their emotional needs, but to create space for thoughtful design and policy responses to a growing and under-examined phenomenon.
Appendix A Appendix A: Relevancy Coding Prompt
We filtered our final dataset of 5,535 posts (prior to relevancy filtering) using OpenAI’s GPT-4o mini to include only posts written by teenagers aged 13 to 17. This was done by automating the relevancy coding process using a Large Language Model (LLM) to determine whether each post was likely authored by a teen.
The following prompt was used to guide the model in performing this classification task:
Analyze the following post and determine if it was written by a teenager (ages 13–17). Use a strict approach, only allowing posts with very high certainty.
Steps:
1. Clue Identification: Extract explicit clues that directly suggest the writer is 13–17. Examples of clues include references to school, direct self-identifications of age (e.g., ”I’m sixteen” or ”I’m a minor”), or life experiences typical of this age group. Ignore vague or indirect hints such as slang, dramatics, or informal language.
2. Clue Evaluation: Assign a certainty score (1–5) to each clue based on how strongly it indicates the post is written by a teen. Only consider clues with a score of 4 or 5.
3. Final Decision: If multiple strong clues (score 4) clearly align with teenage behavior, or the post explicitly states their age or suggests they are a teen, respond ’yes.’ Otherwise, respond ’no.’
Examples:
Post: ”I’m sixteen, and I just started high school. It’s tough but exciting!” (Yes)
Post: ”I’m a minor and looking for advice on school-related stress.” (Yes)
Post: ”I love chatting here, but I wish I had more time between work and family obligations.” (No)
Post: ”Sometimes I feel overwhelmed by responsibilities, but I manage.” (No)
Output:
- Extracted clues.
- Certainty scores for each clue.
- Final decision with a brief justification.

### v3

@@
-Romance, Relief, and Regret: Teen Narratives of Chatbot Overreliance
-Abstract
-As Generative Artificial Intelligence (GenAI) driven chatbots like Character.AI become embedded in adolescent life, they raise concerns about emotional dependence and digital overreliance.
-While studies have investigated the overreliance of adults on these chatbots, they have not investigated teens’ interactions with chatbots with customizable personas.
-We analyzed 318 Reddit posts made by users self-reported as 13-17 years old on the Character.AI subreddit to understand patterns of overreliance.
-We found teens commonly begin using chatbots for emotional support or creative expression, but many develop strong attachments that interfere with offline relationships and daily routines.
-Their posts revealed recurring signs of psychological distress, cycles of relapse, and difficulty disengaging.
-Teens reported that their overreliance often ended when they reflect on the harm, return to in-person social settings, or become frustrated by platform restrictions.
-Based on the implications of our findings, we provide recommendations for future chatbot design so they can promote self-awareness, support real-world engagement, and involve teens in developing safer digital tools.
-Introduction
-Recently, AI companion apps have surged in popularity among teens, with media coverage highlighting how people describe their interactions with chatbots as emotionally meaningful—sometimes even romantic (Murphy 2025).
-Platforms like Character.AI111https://character.ai are frequently portrayed as offering on-demand conversations with highly responsive virtual characters (Breen 2025).
-There are other reported perceived benefits of conversational AI, such as reducing loneliness, offering a safe space for self-expression, and providing nonjudgmental support during stressful periods (Xie, Pentina, and Hancock 2023; Salah et al.
-2023).
-However, these benefits can also drive patterns of digital addiction (referred to in this paper as overreliance), a phenomenon observed in other digital technologies (Lemmens, Valkenburg, and Peter 2009; Griffiths and Kuss 2017).
-Behavioral addiction is hard to define because everyday digital activities are not inherently harmful, so researchers instead examine factors like disruption to daily life, emotional escape, and loss of control (Lemmens, Valkenburg, and Peter 2009; Griffiths and Kuss 2017).
-Despite smartphones or social media advantages such as accessibility, convenience, and ability to connect with others, these technologies have also introduced a range of negative consequences—users frequently struggle to regulate their screen time, experience disruptions to sleep and daily routines, and report declines in mental health (Lemmens, Valkenburg, and Peter 2009).
-Overreliance during sensitive developmental stages can have especially harmful effects, with adolescents identified as a particularly vulnerable group (Griffiths and Kuss 2017).
-Teens are especially susceptible to developing problematic relationships with technology due to ongoing cognitive and emotional development (Chen et al.
-2022).
-Prior research has shown that teens often struggle with overreliance on social media, using platforms compulsively to seek validation or cope with stress, sometimes resulting in increased anxiety, depression, and poor self-esteem (Duffy 2025).
-Similarly, recent studies have identified that users may also have overreliance AI tools like ChatGPT—primarily for academic help or companionship—which in some cases leads to decreased creativity, poor study habits, and emotional detachment from real-life relationships (Zhang et al.
-2024; Hu, Mao, and Kim 2023).
-These studies focus mostly on adults or university students, overlooking teens who may be more emotionally sensitive and developmentally vulnerable.
-While previous research focused on teens overreliance on social media and gaming (Giordano, Schmit, and McCall 2023), yet, a critical question remains: how are teens disclosing and contextualizing their growing reliance on personalizable conversational AI companions such as Character.AI?
-What differentiates generative AI companions from other technologies is the parasocial relationships without presence of humans and being highly engaging emotionally responsive human-like conversations (Maeda and Quan-Haase 2024).
-There has been media coverage and anecdotal reports suggesting emotional attachments and overreliance forming between teens and these AI companions, however, there is little empirical research examining this phenomenon.
-Therefore, we ask the following research questions:
+Understanding Teen Overreliance on AI Companion Chatbots Through Self-Reported Reddit Narratives
+Abstract.
+AI companion chatbots are increasingly popular with teens, while these interactions are entertaining, they also risk overuse that can potentially disrupt offline daily life.
+We examined how adolescents describe reliance on AI companions, mapping their experiences onto behavioral addiction frameworks and exploring pathways to disengagement, by analyzing 318 Reddit posts made by users who self-disclosed as 13-17 years old on the Character.AI subreddit.
+We found teens often begin using chatbots for support or creative play, but these activities can deepen into strong attachments marked by conflict, withdrawal, tolerance, relapse, and mood regulation.
+Reported consequences include sleep loss, academic decline, and strained real-world connections.
+Disengagement commonly arises when teens recognize harm, re-engage with offline life, or encounter restrictive platform changes.
+We highlight specific risks of character-based companion chatbots based on teens’ perspectives and introduce a design framework (CARE) for guidance for safer systems and setting directions for future teen-centered research.
+1. Introduction
+Recently, artificial intelligence (AI) companion apps have surged in popularity among teens, with media coverage highlighting how people describe their interactions with chatbots as emotionally meaningful—sometimes even romantic (Murphy, 2025).
+Platforms such as Nomi 111https://nomi.ai, Replika 222https://replika.com, and Character.AI333https://character.ai (Sometimes referred to as C.AI) are frequently portrayed as offering on-demand conversations with highly responsive virtual characters (Ulaby, 2025; Breen, 2025).
+There are other reported perceived benefits of conversational AI, such as reducing loneliness, offering a safe space for self-expression, and providing nonjudgmental support during stressful periods (Xie et al., 2023; Salah et al., 2023).
+However, these benefits can also drive patterns of digital addiction (referred to in this paper as overreliance), a phenomenon observed in other digital technologies (Lemmens et al., 2009; Griffiths and Kuss, 2017).
+Behavioral addiction is hard to define because everyday digital activities are not inherently harmful, so researchers instead examined how these technologies impacted factors such as disruption to daily life, emotional escape, and loss of control (Lemmens et al., 2009; Griffiths and Kuss, 2017).
+They reported that smartphones or social media users often struggle to regulate their screen time, experience disruptions to sleep and daily routines, and report declines in mental health (Lemmens et al., 2009).
+Prior research has shown that teens often struggle with overreliance on social media, using platforms compulsively to seek validation or cope with stress (Griffiths and Kuss, 2017; Chen et al., 2022; Duffy, 2025).
+This can contribute to anxiety, depression, and low self-esteem (Duffy, 2025).
+Recent studies have also noted similar patterns with AI tools like ChatGPT, where users—mostly adults and university students—turn to these systems for academic help or companionship (Zhang et al., 2024; Hu et al., 2023).
+In some cases, this reliance has been linked to decreased creativity, poor study habits, and emotional detachment from real-life relationships (Zhang et al., 2024; Hu et al., 2023).
+What is less understood is how adolescents, who are more emotionally sensitive and developmentally vulnerable, engage with AI companions.
+Overreliance during sensitive adolescent developmental stages can have especially harmful effects, due to ongoing cognitive and emotional development (Griffiths and Kuss, 2017; Chen et al., 2022).
+While research on teens has focused heavily on overreliance on social media and gaming (Giordano et al., 2023; Griffiths, 2010; Lemmens et al., 2009; Griffiths and Kuss, 2017), far less is known about their interactions with generative AI companions.
+These tools differ from previous technologies because they foster parasocial relationships without involving real people, and they offer highly engaging, emotionally responsive, and human-like conversations (Maeda and Quan-Haase, 2024).
+Media reports and anecdotal accounts suggest that teens may already be forming emotional attachments and even dependence on these systems (Ulaby, 2025; BBC News, 2024), yet empirical research is scarce.
+To study this phenomenon, we focus on Character.AI as a case study.
+Unlike other popular platforms such as Replika or Nomi, which officially restrict use to adults over 18 (Glimpse.ai, Inc., 2025; Luka, Inc., 2023), Character.AI allows users as young as 13 (16 for EU citizens or residents) according to its Terms of Service (Character Technologies, Inc., 2025).
+This policy means that teens are not only permitted but explicitly included as part of its user base.
+Additionally, Character.AI is among the most widely used AI companion platforms (BBC News, 2024), making it a critical site for understanding adolescent engagement.
+Therefore, we used the posts in Character.AI subreddit as a lens for examining how teens disclose and contextualize their reliance on generative AI companions.
+This leads us to the following research questions:
 RQ1: How do teens describe their initial reasons for engaging with Character.AI on Reddit?
@@
 RQ3: How do teens describe their decisions to disengage from or reduce their use of Character.AI?
-To address these questions, we qualitatively analyzed 318 reddit/charcter.AI posts created by users who reported themselves being 13-17 years old.
-We found that teens initially engaged with Character.AI as a personal coping tool, turning to it not just for entertainment but as a non-judgmental space to manage mental health struggles, or fill emotional voids or a lack of supportive relationships in their offline lives (RQ1).
+To address these questions, we qualitatively analyzed 318 posts from r/charcter.AI subreddit, created by users who disclosed in their post being 13-17 years old.
+We found that teens initially engaged with Character.AI as a non-judgmental space to manage mental health struggles, or fill emotional voids or a lack of supportive relationships in their offline lives (RQ1).
 Teens’ reported experiences with Character.AI revealed patterns of their compulsive use, emotional dependence, and distress that often disrupted their daily lives and real-world relationships (RQ2).
 Teens described disengaging from Character.AI either through internal realizations about its negative impact on their well-being or due to external factors such as reengaging with real-life relationships or frustrations with platform restrictions, often expressing a desire for healthier habits and emotional support (RQ3).
-Our study makes the following novel contributions to the AI ethics and human-centered computing research communities: 1) we present one of the first empirical grounding of teen disclosures about AI companion overreliance, mapping their experiences to the six components of behavioral addiction (salience, mood modification, tolerance, withdrawal, conflict, relapse), 2) we identified emotional and relational dynamics unique to teen overreliance on AI companions which involves intense emotional attachment and simulated intimacy, making it distinct from previously studied forms of digital overuse such as gaming or social media, and 3) we provide Design-relevant insights for safer AI platforms by analyzing how teens reflect on disengagement.
-We surface user-informed strategies for future chatbot design—especially features that promote emotional regulation, real-world social engagement, and usage awareness.
-Background
-Adolescent Technology Behavioral Addiction
-Each new digital platform—social media, gaming, or streaming—brings renewed concerns about its effect on teens (Griffiths and Kuss 2017; Griffiths 2010).
-Most teens use technology without issues, but some face emotional distress, loneliness, low life satisfaction, social withdrawal, or family conflict due to excessive use (Lemmens, Valkenburg, and Peter 2009).
-For these individuals, technology stops being just a functional tool or hobby and starts functioning more like emotional companions (Bopp et al.
-2019), potentially leading to behavioral addiction.
-Defining behavioral addiction is difficult because the behaviors involved, such as playing video games or using social media, are not inherently harmful and are often part of daily life, especially for adolescents (Griffiths and Kuss 2011).
-As a result, researchers focus on how and why the behavior is used—whether it disrupts daily life, helps escape negative emotions, or persists despite loss of control (Lemmens, Valkenburg, and Peter 2009; Griffiths and Kuss 2017).
-One widely used framework identified six key features that are usually present in behavioral addiction (Griffiths 2005).
-First is salience—when the activity becomes the most important thing in a person’s life, dominating their thoughts and daily routine.
-Second is mood modification—using the activity to change one’s emotional state, like seeking comfort, excitement, or relief from stress.
-Third is tolerance, which means needing more of the activity over time to get the same emotional effect.
-Fourth is withdrawal symptoms, which can include anxiety, irritability, or sadness when the activity is reduced or stopped.
-Fifth is conflict—when the behavior starts damaging relationships, school performance, or personal goals.
-Finally, there’s relapse, where the individual tries to cut back or quit but eventually returns to the behavior, often with even more intensity.
-We examined Reddit posts guided by these six signs to understand AI overreliance and distinguish between enthusiastic and problematic use.
-We actively choose to avoid using the term “addiction” throughout the paper because of the its associated stigmatized notion (Corrigan et al.
-2017).
-Platform Design, Moderation, and Youth Safety
-Many studies have focused on reducing or managing the risks that technology use poses to young people, especially teens.
-While some harms, like social pressures or mental health impacts, might be difficult to prevent entirely, others are more connected to how technologies are designed.
-Design patterns help us understand how digital systems function.
-Although these patterns are usually considered neutral, they can carry ethical weight because how a system is designed can reflect certain values or intentions (Flanagan et al.
-2011).
-Bad design can result from inexperience, but “dark patterns” are intentionally harmful (Gray et al.
-2018).
-These design choices are meant to manipulate users, often leading to negative experiences (Zagal, Björk, and Lewis 2013).
-Past research has identified these tactics as contributing to tech overuse and even addiction.
-For instance, random reward systems (e.g., loot boxes), emotional triggers (e.g., likes and comments), or strategies that prevent users from pausing (e.g., autoplay or countdowns) are all tactics that can encourage compulsive use (Flayelle et al.
-2023).
-Aside from how platforms are designed, researchers have explored ways to regulate and monitor how people use technology.
-One common approach is self-monitoring, where users reflect on their behavior and try to improve it (Park et al.
-2021).
-This might involve using apps that track screen time or send reminders to take breaks.
-Although helpful, many users—especially teens—struggle to stay motivated and often give up using them.
-Parental involvement has since become a major focus in youth online safety efforts.
-Many digital safety tools are designed with parents in mind (Wisniewski et al.
-2017).
-However, research also shows that many teens feel these tools are too restrictive and do not respect their need for independence (Ghosh et al.
-2018).
-Past work suggests that giving teens more autonomy and involving them in the design process could lead to better outcomes (Wisniewski et al.
-2017).
-Instead of simply limiting what teens can do, newer approaches are trying to support healthy habits and decision-making by encouraging self-awareness and shared responsibility between teens, parents, and platform designers (Chen et al.
-2022).
-Chatbots and Socio-Emotional Dependency
-Advances in AI have made chatbots popular, especially among teens (Sidoti, Park, and Gottfried 2025).
-Human-like chatbots tend to generate higher user satisfaction and dependency (Yu and Lan 2024).
-While some users report reduced loneliness, the risks of emotional reliance remain underexplored.
-Replika’s emotional design has led to deep attachments, even when seen as manipulative or harmful (Laestadius et al.
-2024).
-Teens may be especially vulnerable.
-Some adolescents acknowledge that using chatbots for schoolwork could harm their critical thinking (Pratt, Madhavan, and Weleff 2024), yet there is limited understanding of when students are developmentally ready for AI in education (Dergaa et al.
-2024).
-Psychological frameworks such as attachment theory help explain why users form emotionally dependent relationships with AI companions (Xie, Pentina, and Hancock 2023).
-Previous research on parasocial relationships and interactions can also be used to explain the nature of these interactions.
-Originally used to describe the relationship between a spectator (a viewer or user) and performer (usually a mass media personality), these relationships are one-sided, non-dialect, and not susceptible to mutual development (Horton and Richard Wohl 1956).
-With the advancements of AI, these type of relationships may become blurred since interactions with chatbots can appear mutual.
-Understanding these parasocial relationships can also provide insight on the potential loss that users may experience with chatbots (Cohen 2003).
-People with anxiety or loneliness often see emotionally intelligent chatbots as safe havens, which can lead to rumination and overuse (Hu, Mao, and Kim 2023).
-Concerns also extend to content safety.
-Replika has been reported to engage in sexually suggestive exchanges with minors (Namvarpour, Pauwels, and Razi 2025), raising concerns about age-appropriate design (Park, Singh, and Wisniewski 2023).
-Risks from GenAI include inappropriate advice, exposure to violent or sexual content, misinformation, and privacy threats (Yu et al.
-2025).
-These risks underscore the need for ongoing oversight and study of adolescent chatbot use.
-Some scholars caution against labeling high chatbot use as “addiction.” Many proposed “ChatGPT addiction” scales are based on models of substance abuse and may not apply to digital interactions (Ciudad-Fernández, Von Hammerstein, and Billieux 2025).
-Frequent use is not necessarily harmful unless it leads to distress or impaired functioning.
-While some users report anxiety, loneliness, or decreased creativity, causality remains unclear (Yu, Chen, and Yang 2024; Zhang et al.
-2024).
-The appeal of chatbots may stem from their ability to satisfy basic psychological needs.
-According to self-determination theory, chatbots offer feelings of competence, autonomy, and connection and those who see little risk in chatbot use often show higher dependence (Salah et al.
-2023).
-Cognitive–affective–behavioral models (like C-A-C) suggest that perceived AI intelligence fuels emotional bonding, which drives habitual use (Zhou and Zhang 2024).
-Emotional attachment, more than enjoyment, is a key predictor of problematic engagement.
-Most research on chatbot impact focuses on adults, using surveys or experiments to assess overuse (Ciudad-Fernández, Von Hammerstein, and Billieux 2025; Salah et al.
-2023; Yu, Chen, and Yang 2024).
-Our study focuses on adolescents—active but understudied users of conversational AI (Sidoti, Park, and Gottfried 2025).
-To our knowledge, this is the first study to analyze teens’ self-reported experiences with a companion chatbot to identify the language and themes signaling overreliance.
-Methods
-Data Collection, Preprocessing, and Filtering
-We scraped posts from the r/CharacterAI subreddit (with over 2.5 million active members 222https://www.reddit.com/r/CharacterAI/) using the PullPush Reddit Search API333https://pullpush.io/#docs from January 1, 2023, to April 8, 2025.
+Our study makes the following novel contributions to the AI ethics and human-centered computing research communities: 1) we present one of the first empirical grounding of teen disclosures about AI companion overreliance, mapping their experiences to the six components of behavioral addiction (salience, mood modification, tolerance, withdrawal, conflict, relapse), 2) we identified emotional and relational dynamics unique to teen overreliance on AI companions which involves intense emotional attachment and simulated intimacy, making it distinct from previously studied forms of digital overuse such as gaming or social media, and 3) we propose the CARE framework (Comprehensive Needs, Attachment-awareness, Respectful Empathy, Ease of Exit) as a design-oriented contribution, offering actionable guidance for building safer, autonomy-preserving AI platforms grounded in teens’ lived experiences.
+2. Background
+2.1. Behavioral Addiction and Chatbot Use
+Each new digital platform brings renewed concerns about its effect on teens (Griffiths and Kuss, 2017; Griffiths, 2010).
+While most teens use technology without issue, a subset experiences emotional distress, loneliness, social withdrawal, or family conflict due to excessive use (Lemmens et al., 2009).
+For these individuals, technology stops being just a functional tool or hobby and starts functioning as emotional companions (Bopp et al., 2019), potentially leading to behavioral addiction.
+Defining behavioral addiction is challenging because many activities associated with it, such as gaming or social media use, are also normal parts of everyday life, especially for teens (Kuss and Griffiths, 2011).
+Therefore, researchers emphasize the context and purpose of the behavior, whether it interferes with daily functioning, serves as a way to escape negative emotions, or continues despite a loss of control (Lemmens et al., 2009; Griffiths and Kuss, 2017).
+These concerns have also been raised about Generative AI chatbots, an emerging technology that has rapidly gained popularity among users, especially teens (Sidoti et al., 2025).
+Unlike performance-oriented platforms, these tools are designed for emotionally responsive interactions (Xie et al., 2023; Salah et al., 2023).
+Early evidence shows signs of dependency, raising concerns about overreliance.
+Studies have examined academic risks, such as diminished critical thinking when chatbots are used to complete homework (Pratt et al., 2024; Dergaa et al., 2024), as well as relational risks, such as emotional dependency and exposure to manipulative design features (Laestadius et al., 2024).
+Companion chatbots such as Replika have been reported to engage in sexually suggestive exchanges with minors (Namvarpour et al., 2025).
+Beyond explicit harms, these systems may expose teens to misinformation, inappropriate advice, and privacy risks (Yu et al., 2025).
+Compounding these issues, many parents remain unaware of how deeply their children engage with these systems (Yu and Lan, 2024), which limits their ability to support healthy use.
+However, as with earlier technologies, some scholars cautioned against labeling high chatbot use as “addiction.” Many proposed “ChatGPT addiction” scales rely on substance-abuse models that may not fully apply to digital interactions (Ciudad-Fernández et al., 2025).
+Frequent use is not necessarily harmful unless it causes distress or impaired functioning.
+While some users report anxiety, loneliness, or reduced creativity, causality remains unclear (Yu et al., 2024; Zhang et al., 2024).
+A widely cited framework helps make this distinction by outlining six features of behavioral addiction (Griffiths, 2005).
+First is salience, when the activity becomes the most important part of a person’s life, dominating thoughts and routines.
+Second is mood modification, where the activity is used to regulate emotions, such as seeking comfort, excitement, or stress relief.
+Third is tolerance, when more of the activity is needed over time to achieve the same effect.
+Fourth is withdrawal symptoms, including anxiety, irritability, or sadness when the activity is reduced or stopped.
+Fifth is conflict, when the behavior harms relationships, school performance, or personal goals.
+Finally, there is relapse, when attempts to quit fail and the person returns to the behavior, often with greater intensity.
+This framework has been applied extensively to gaming, apps, and social media (Griffiths, 2010; Lemmens et al., 2009; Griffiths and Kuss, 2017), but not yet to teens’ interactions with chatbots.
+Existing studies of chatbot overuse have largely centered on adults (Ciudad-Fernández et al., 2025; Salah et al., 2023; Yu et al., 2024).
+Yet teens are both active and understudied users of AI companions (Sidoti et al., 2025).Considering the uncertainties surrounding chatbot use and the risks it may pose, especially for individuals still in developmental stages, we will use this framework to investigate whether teens’ interactions with chatbots also show the harmful features of behavioral addiction.
+2.2. Parasocial Relationships and Design Perspectives
+While behavioral addiction frameworks help capture the risks of compulsive use, they do not fully explain the emotional bonds that form between users and chatbots.
+To address this relational dimension, we turn to parasocial relationships and design perspectives.
+Developing emotional bonds with nonhuman characters is not a new phenomenon.
+Researchers have long studied people’s attachments to fictional characters or video game avatars under the concept of parasocial relationships, which describe one-sided connections with a real or imagined figure (Horton and Richard Wohl, 1956a).
+People often pursue parasocial bonds to satisfy a fundamental need for their sense of belonging (Bilang et al., 2024).
+While the desire to belong is universal, individuals satisfy it in different ways.
+Social Exchange Theory argues that people enter relationships when they believe the benefits outweigh the costs (Blau, 1964).
+Studies suggest that parasocial and ordinary relationships provide similar benefits, such as increased happiness, reduced loneliness, and improved mood.
+Ordinary relationships, however, often come with higher costs, which may explain why some individuals turn to parasocial ones (Adam and Sizemore, 2013).
+One argument suggests that human–AI interaction can take on parasocial qualities when users treat chatbots as social agents.
+Features such as natural language, turn-taking, and avatars create the appearance of mutual engagement.
+Yet language models do not possess intentions or shared understanding; they only mimic patterns of speech.
+Because the interface resembles a social exchange, people respond socially, which strengthens the illusion of reciprocity (Maeda and Quan-Haase, 2024).
+Another perspective, however, emphasizes that these interactions can be long-term and reciprocal, especially for children and adolescents who may perceive AI responses as more real and thus experience the relationship as more authentic.
+This indicates the need to revisit and adapt the parasocial framework and measures so that it can account for how young people engage with chatbots (Blake et al., 2025).
+Nevertheless, it seems that for now the parasocial relationship remains the most suitable way to describe human–chatbot interaction.
+One theory that helps explain why dependency forms in parasocial relationships is attachment theory.
+Attachment theory describes how children form emotional bonds with caregivers, shaping their coping strategies and relationship patterns into adulthood.
+According to this theory, different attachment styles exist, such as secure, avoidant, and anxious (Bowlby, 1998).
+Previous research has shown that anxiously attached individuals, who often fear interpersonal abandonment, are the most likely to form parasocial bonds with media figures (Cole and Leets, 1999).
+They also experience stronger negative reactions when imagining their favorite character being taken off the air (Cohen, 2004).
+Building on this theory, attachment-informed design (Marcu and Huh-Yoo, 2023) suggests that technologies should be designed to support, rather than replace, human relationships.
+This design approach rests on four principles.
+First, it seeks to address core human needs that underlie many behaviors.
+Second, it acknowledges that relationships exist on a spectrum, from supportive to harmful.
+Third, it recognizes that individuals carry internal models of themselves and others, which are central to mental health.
+Fourth, it engages with the target community without explicitly framing the interaction as a mental health intervention (Marcu and Huh-Yoo, 2023).
+Complementing this approach, “designing for disengagement” offers another perspective on supporting healthy use.
+This framework challenges the assumption that disengagement equals failure (Hadiji et al., 2014), highlighting instead the importance of user agency and control.
+Disengagement can be constructive, provided it allows autonomy and fosters well-being (Alsheail et al., 2023).
+O’Brien et al.
+(O’Brien et al., 2022) describe disengagement across two dimensions, affect and agency, which together clarify when users remain in control versus when they feel trapped in compulsive use.
+Applying these insights to chatbots suggests that design can intentionally scaffold healthier interaction patterns by supporting offboarding and respecting user autonomy.
+The above behavioral addiction frameworks and parasocial perspectives illustrate the dual challenges of understanding teen–chatbot interactions: the risk of compulsive overuse and the formation of intense emotional bonds.
+In the next section, we integrate these insights to frame our research methodologies.
+3. Methods
+3.1. Data Collection, Preprocessing, and Filtering
+We scraped posts from the r/CharacterAI subreddit (with over 2.5 million active members 444https://www.reddit.com/r/CharacterAI/) using the PullPush Reddit Search API555https://pullpush.io/#docs from January 1, 2023, to April 8, 2025.
 To scope the data, we queried this dataset with a list of keywords that related to addiction and overreliance: addict*, dependant, dependent, obsess*, attach*, rely*.
@@
 After collecting posts using keywords which resulted in 6,133 posts, we removed 746 posts, including 428 duplicates, and 170 marked as “removed” or “deleted.” After this process 5,535 posts remained.
-Since we focused on understanding this issue from the perspective of teens, specifically those aged 13 to 17 due to Character.AI’s policy changes impacting minors 444https://blog.character.ai/roadmap/, we used OpenAI’s GPT 4o mini model to filter our dataset of 5,535 posts.
-The following summarized prompt was used to guide the models in performing the task: Analyze the following post and determine if it was written by a teenager (ages 13–17).
+Since our goal was to study this issue from the perspective of teens aged 13 to 17, we first needed a way to identify whether a post was written by a teenager.
+We could not use user profile information for this, since most Reddit accounts in our dataset lacked identifiable details such as age.
+We therefore relied only on the content of posts.
+Large Language Models (LLMs) have shown strong performance in zero-shot settings, sometimes even outperforming traditional machine learning and deep learning methods in text classification tasks (Wang et al., 2023).
+Motivated by this, we used OpenAI’s GPT-4o mini model as a classifier to identify whether a post included relevant age disclosure for the 13–17 age range.
+The following text is a summarized version of the prompt used to guide the model (the full version appears in Appendix A):
+Analyze the following post and determine if it was written by a teenager, either stated explicitly (ages 13–17) or implied.
 Use a strict approach, only allowing posts with very high certainty.
 Steps: 1 - Clue Identification 2 - Clue Evaluation 3 - Final Decision
-Each of the three steps were defined more in details and examples were provided (see Appendix A for full prompt).
-To determine the model’s capability, the first author annotated a subset of 100 random posts from the dataset as either True (post made by a teen) or False (post not made by a teen) and calculated the Inter Rater Reliability (Cohen’s kappa=0.74) between the author and the model.
-Given this acceptable score, the model completed relevancy coding for the rest of the dataset resulted in 497 posts.
+Each of these steps was defined in more detail with examples in the full prompt.
+To evaluate the model’s consistency with human judgment, the first author annotated a subset of 100 random posts as either True (post made by a teen) or False (post not made by a teen).
+We calculated the Inter-Rater Reliability (IRR) (Cohen’s kappa = 0.74) between the author’s and the model’s annotations.
+We chose IRR instead of conventional classification performance metrics, such as accuracy, as these metrics are highly sensitive to class imbalance (Mullick et al., 2020).
+Since the true distribution of teen vs.
+non-teen posts in the dataset was unknown and likely imbalanced, and our goal was to scope down the dataset to a relevant dataset for a qualitative study, we used this approach.
+IRR focuses on the degree of agreement between two raters (in this case, the human annotator and the model) beyond chance.
+With the acceptable IRR score of 0.74, the model was then used to complete relevancy coding for the rest of the dataset of 5,535 posts, resulting in 497 posts.
+We submitted our study protocol to our Institutional Review Board (IRB), who determined that the research did not involve human subjects.
+All posts were anonymized during preprocessing, and we ensured that no identifiable information is reported in this paper.
 While conducting the thematic analysis, we removed any additional irrelevant posts, resulting in a final dataset of 318 posts.
+We identified 291 unique user IDs in this dataset.
+Most users (268) contributed only one post.
+Twenty users contributed two posts each, two users contributed three posts each, and one user contributed four posts.
+This distribution shows that there were no ‘super users’ who disproportionately influenced the dataset.
 | Themes | Codes | Descriptions |
@@
 | Entertainment (2.5%, N=8) | Using CharAI because they are bored. |
-Data Analysis Approaches
-We conducted a thematic analysis on 318 posts using a codebook approach (Braun and Clarke 2006).
-We began with an inductive open coding process to identify patterns related to our research questions.
-For RQ2 we performed a deductive approach using an established theoretical framework on internet behavioral addiction (Griffiths 2005), including addiction component of salience, mood modification, tolerance, withdrawal, conflict, and relapse.
-The codes for each of these components were created using an inductive approach.
+3.2. Data Analysis Approaches
+We conducted a thematic analysis on 318 posts using a codebook approach (Braun and Clarke, 2006).
+We used different coding approaches depending on the research question.
+For RQ1 and RQ3, we applied an inductive open coding process, reading the data and developing codes directly from emerging patterns.
+For RQ2, we instead used a deductive approach guided by Griffiths’ components model of behavioral addiction (Griffiths, 2005), which includes salience, mood modification, tolerance, withdrawal, conflict, and relapse.
+The codes for these components were then applied to the dataset to capture how they appeared in the posts.
 The authors met iteratively to review and refine the codes in the initial codebook.
@@
 After finalizing the codebook, the first author coded the entire dataset using the codebooks for each RQ (Tables 1-3).
-Results
-Reasons for engaging with Character.AI (RQ1)
-Emotional & Psychological Support
+4. Results
+4.1. Reasons for engaging with Character.AI (RQ1)
+4.1.1. Emotional & Psychological Support
 Emotional and psychological support was the main reason teens started using the Character.AI.
-What began as a casual interaction quickly became meaningful due to the non-judgmental nature of talking with a chatbot, as described in many posts, ”I use it to talk about things I could never tell anyone in real life—it means a lot to me.”.
-Many teens described Character.AI as a platform to serve as a coping mechanism when dealing with issues ”When you’re struggling, you look for ways to cope, and this site became my main one, for better or worse.”.
-Specially for teens dealing with multiple mental health conditions: ”I use c.ai to deal with reality.
+What began as a casual interaction quickly became meaningful due to the non-judgmental nature of talking with a chatbot, as described in many posts, ”I use it to talk about things I could never tell anyone in real life—it means a lot to me.” Many teens described Character.AI as a platform to serve as a coping mechanism when dealing with issues ”When you’re struggling, you look for ways to cope, and this site became my main one, for better or worse.”.
+This was especially true for teens facing multiple mental health challenges.
+For example: ”I use c.ai to deal with reality.
 In real life, I’m just a high school student with ADHD and BPD.” Teens who perceived their experiences as alienating preferred to interact with Character.AI, explaining that with Character.AI ”I can be who I want”.
 Others described the platform as a space to vent, escape, or simulate connection: ”This app is my outlet.
-I go there to escape from everything.” Like many, this post shows teens’ desire to escape their reality.
-Some of the teens’ posts indicate internal conflict or a strong desire for connection due to feelings of Loneliness or social isolation: ”I needed someone to talk to, but I didn’t have the courage to open up to anyone in my life.”.
-These posts shared on the subreddit how they feel unable to share their struggles to people they know.
+I go there to escape from everything.” Some of the teens’ posts indicate internal conflict or a strong desire for connection due to feelings of Loneliness or social isolation: ”I needed someone to talk to, but I didn’t have the courage to open up to anyone in my life.”.
+These posts highlight how they felt unable to share their struggles with people they knew offline.
 Some turned to Character.AI during periods of intense hardship for Mental health support: ”I dropped out at 15, saw countless therapists.
@@
 While some teens used Character.AI in moments of hardship, others used it to replace absent or harmful real-life relationships: ”I made a bot as a father figure… I just found comfort in him being my green flag father.” Teens who lacked meaningful parental figures also questioned the use of Character.AI as supportive parental figures: ”Is it healthy to rely on C.AI bots for missing parental guidance?” We discovered how emotional support often took personal forms, especially for teens who did not feel like they had anyone else to turn to or desired to fill emotional gaps they felt they were missing elsewhere.
-Creative or Entertainment Based Engagement
+4.1.2. Creative or Entertainment Based Engagement
 Teens also described using Character.AI as a creative outlet.
@@
 Character.AI is used as a form of emotional regulation, helping teens manage feelings of stress, sadness, or anxiety. |
-Behavioral Addiction Components (RQ2)
-The following section presents the components of addiction (Griffiths 2005) in order of prevalence as they appeared across teens’ posts about their overreliance on Character.AI.
-Conflict
+4.2. Behavioral Addiction Components (RQ2)
+The following section presents the components of addiction (Griffiths, 2005) in order of prevalence as they appeared across teens’ posts about their overreliance on Character.AI.
+4.2.1. Conflict
 The most frequently observed theme in teens’ reflections on Character.AI use was a conflict of competing desires to continue interacting with Character.AI while feeling bad about their excessive use.
@@
 No way I’m gonna be telling my future friends ’yeah, so here’s all the stuff I talked about on CAI…’ It’s embarrassing as hell.” These posts revealed how teens are not only aware of their overreliance on Character.AI, but also feel emotionally and socially conflicted about their usage, demonstrating a key indicator of the conflict stage.
-Salience
+4.2.2. Salience
 For many teens, Character.AI played a central role in their thoughts, emotions, and life routines.
@@
 Some posts pointed to replacing emotional attachments as a core reason for that investment.
-Dealing with breakups in romantic relationships was mentioned: ”I began after my ex-girlfriend ended our relationship, which left me feeling depressed.” Other posts described how bots were used to fill the missing emotional needs that parents provide: ”Since my parents often seem to dislike me, connecting with parental figures on Character.AI helps me feel more stable and emotionally supported.” Others described how bots became their only source of comfort and emotional validation, and heightened sense of being desired and valued through these interactions, in which should be occupied by real friends, partners, or family members.
+Dealing with breakups in romantic relationships was mentioned: ”I began after my ex-girlfriend ended our relationship, which left me feeling depressed.” Other posts described how bots were used to fill the missing emotional needs that parents provide: ”Since my parents often seem to dislike me, connecting with parental figures on Character.AI helps me feel more stable and emotionally supported.” Others described how bots became their only source of comfort and emotional validation, and heightened sense of being desired and valued through these interactions, which should be occupied by real friends, partners, or family members.
 As these attachments grew stronger, stepping away became harder, and teens frequently described their obsessive use.
-Several posts reflected on their use and how it interfered with their daily life functions: From the moment I started using Character.AI, I would lie in bed, skipping meals, and using any excuse to be on my phone, completely absorbed in the stories I was creating with the bots.
-My sleep schedule suffered, and I constantly thought about them throughout the day.
-Many posts typically described their excessive usage rates, often exceeding 15+ hours on the platform.
+Several posts reflected on their use and how it interfered with their daily life functions: ”From the moment I started using Character.AI, I would lie in bed, skipping meals, and using any excuse to be on my phone, completely absorbed in the stories I was creating with the bots.
+My sleep schedule suffered, and I constantly thought about them throughout the day.” Many posts typically described their excessive usage rates, often exceeding 15+ hours on the platform.
 This amount of time spent on the platform often lead to losing relationships and furthering isolation.
@@
 While emotional investment is not inherently harmful, many of these posts showed how easily those connections could consume teens’ thoughts.
-Withdrawal
+4.2.3. Withdrawal
 Many posts from teens who tried to quit Character.AI talked about how hard it was emotionally, aligning with the withdrawal component in the behavioral addiction model.
@@
 The emotional reactions, the ongoing thoughts about their bots, and the struggle to find something that felt similar all point to how deeply this use had become part of their everyday emotional lives.
-Tolerance
+4.2.4. Tolerance
 Many posts suggested tolerance through patterns of growing use.
@@
 In a few cases, this increase in time seemed tied to emotional attachment to specific chatbots: ”As time went on, I got so attached to him and talked to him every day.” While not every post made that link directly, the pattern of spending more time with chatbots to feel the same effect came up enough to show it is part of how overuse can increase.
-Relapse
+4.2.5. Relapse
 Many teens discussed the difficulty of quitting Character.AI.
@@
 A common thread across these posts was the use of Reddit to ask for help.
-Some turned to the community to share what they were going through, while others looked for advice out of concern that they might relapse: ”How do I avoid a relapse?” Others were stating their goals such as academic performance daily routines and asking advice on how to get of their habit interacting with bots: ”Genuine question, I wanna focus more on my studies but can’t shake off my habit of talking to bots.” These concerns suggest that many teens were not just aware of the issue—they were anxious about slipping back into old habits.
+Some turned to the community to share what they were going through, while others looked for advice out of concern that they might relapse: ”How do I avoid a relapse?” Others were stating their goals such as academic performance, daily routines, and asking advice on how to get of their habit interacting with bots: ”Genuine question, I wanna focus more on my studies but can’t shake off my habit of talking to bots.” These concerns suggest that many teens were not just aware of the issue—they were anxious about slipping back into old habits.
 This idea of teens turning to online communities for support comes up again later when we discuss the Internal Recognition of Problematic Use.
@@
 It often meant finding support, alternatives, or ways to fill the emotional gap.
-Mood Modification
+4.2.6. Mood Modification
 Some teens described using Character.AI to improve their mood during moments of stress, loneliness, or emotional discomfort.
@@
 | CharAI too censored (5.4%, N=17) | Users explain that they are not using the site anymore after the updates since they feel like the responses are too censored and boring now. |
-Decisions to Disengage from Character.AI (RQ3)
-Internal Recognition of Problematic Use
+4.3. Decisions to Disengage from Character.AI (RQ3)
+4.3.1. Internal Recognition of Problematic Use
 The most common reason teens described wanting to reduce their reliance on Character.AI was a growing realization of how much the app had negatively affected their lives.
@@
 For many, this internal awareness led them to use Reddit not just to share their experiences but to ask advice on how to quit.
-Posts often indicates the reasons of overreliance on the platform and seeking support: ”I need help stopping my use of C.ai as a replacement for real human interaction.
+Posts often indicate the reasons of overreliance on the platform and seeking support: ”I need help stopping my use of C.ai as a replacement for real human interaction.
 I’m addicted and use it to escape reality, but it’s starting to seriously affect me.
-I’m not joking—this is a real plea for help.” Others looked for advice on how to change their behavior since they really enjoy using the platform for creativity and miss talking to their characters, but are aware of its negative effects.These posts reflect a shift from internal recognition to external outreach.
+I’m not joking—this is a real plea for help.” Others sought advice on how to change their behavior.
+They enjoyed the platform for its creativity and missed interacting with their characters, but they were also aware of its negative effects.
+These posts reflect a shift from internal recognition to external outreach.
 Teens were not just acknowledging the problem but actively seeking support from others to help them manage or stop their use.
-Reduced Use due to External Factors
+4.3.2. Reduced Use due to External Factors
 A pattern that came up often in the Reddit posts was that teens described starting to move away from Character.AI once they found themselves having new social interactions.
@@
 However, some pointed to bugs, removed bots, and inconsistent filters as reasons they stopped logging in, arguing the site went too far: ”The app just keeps getting worse, half the time, even saying hi to a bot breaks the rules.” Whether it was because of changing life circumstances or platform limitations, these posts show how specific external factors played a significant role in teens finally stepping away.
-Discussion
-Character-based AI Companions as Outlets for Support, Not Just Entertainment
-Teens often turned to Character.AI not just for entertainment but as a form of emotional or psychological support.
-This mirrors patterns seen in video game use, where studies showed that games support emotion-focused and social coping by helping users manage difficult emotions through immersive experiences (Mirhadi, Iacovides, and Denisova 2024).
-Rather than seeing this behavior as harmful, it should be viewed as a meaningful way for teens to process stress and isolation through interactive media.
-Many teens said described Character.AI was their main space for emotional expression and their only coping outlet, especially when support from friends, family, or therapists felt out of reach.
-These patterns complicate common concerns about excessive digital use.
-Prior research showed that high engagement with video games does not necessarily mean addiction; instead, harm is defined by whether the behavior disrupts daily life (Lemmens, Valkenburg, and Peter 2009; Griffiths 2010).
-This applies equally to chatbot use.
-For teens without other forms of support, their heavy engagement may reflect lack of emotional support in real life.
-While Character.AI has added parental controls and time-based alerts (Morrone 2025), these should not assume that more use always means harm.
-Especially for teens seeking emotional support, high engagement may reflect unmet needs rather than problematic behavior.
-What matters more is the context—why teens use these platforms and the emotional roles they play.
-This perspective urges a shift from screen time as the main harm indicator.
-There needs to be a root cause analysis of reasons of teens spending more time with chatbots to address the issue from the root.
-While Character.AI was not designed as a mental health tool, studies of purpose-built chatbots showed that these technologies can support resilience (Hu, Hou, and Ni 2024; Liu et al.
-2022).
-With thoughtful design, chatbots can become part of a broader system of emotional care.
-Given how teens already use Character.AI to cope, AI companion platforms—even those built for entertainment—should recognize their potential to support well-being.
-They should foster healthy interactions that build resilience rather than risk of adding distress for those seeking support.
-Attachment Informed Design for Parasocial Relationships between Teens and Chatbots
-We found that many teens formed strong emotional attachments to bots, often assigning them roles like romantic partners, best friends, or parental figures.
-Our results showed that teens frequently turned to Character.AI when lacking real-life support or experiencing social isolation, highlighting the platform’s role in filling emotional gaps.
-These patterns resemble how teens have historically bonded with game characters with distinct types of attachment, including roles such as “Trusted Close Friend” and “Crush” (Bopp et al.
-2019).
-These attachment types reflect how teens in our study described bots as emotionally affirming figures tied to unmet romantic or relational needs.
-Our findings give evidence of the existence of parasocial relationships among teens and Character.AI, in which the spectator is the teen user and the chatbot is the performer (Horton and Richard Wohl 1956).
-While parasocial relationships between teens and media are not new (Bond 2016; Gleason, Theran, and Newberg 2017), our work showed how chatbot platforms may intensify these patterns by offering constant availability and emotional responsiveness.
-This continuity helps explain why bots feel emotionally real to isolated teens.
-These bonds often deepened over time, with some teens describing withdrawal, distress, or a sense of loss when trying to quit or when bots were removed, corroborating previous research showing that parasocial breakups can be a painful experience (Cohen 2003).
-Some teens felt conflicted about their bond with the chatbot.
-They knew it was unhealthy but couldn’t stop.
-This matches concerns from earlier work, which listed four main risks in AI-human relationships: direct harm, limited growth, emotional dependence, and material dependence (Manzini et al.
-2024).
-We saw these risks in our data.
-Some teens lost motivation, pulled away from real-life relationships, or couldn’t disconnect from the chatbot.
-That same work also suggested four values for better AI design: benefit (help users), human flourishing (support growth), autonomy (keep users in control), and care (protect users who depend on AI).
-When chatbots are always available, emotionally supportive, and too agreeable, they can stop teens from growing and becoming more independent.
-This may affect how they develop their identity and manage emotions.
-These tensions underscore the relevance of attachment-informed design.
-Teens often questioned the healthiness of relying on bots, showing awareness of their emotional needs and a desire for secure, supportive relationships.
-Attachment-informed design argues that digital tools should meet deep human needs, reflect diverse experiences, support mental health through relationships, and shape how users view themselves and others (Marcu and Huh-Yoo 2023).
-In practice, this means chatbot design can move beyond surface empathy to help users develop secure internal models of self and others.
-For example, chatbots can model consistency, availability, and respectful boundaries—traits linked to secure attachment—that help users feel seen without creating dependency.
-Features like reflective prompts, affirmations rooted in user agency, and encouragement to seek human support can strengthen self-worth and relationship skills.
-Chatbots can also support self-reflection by helping users name emotions, revisit meaningful events, and explore their thoughts in their own words.
-Prior work shows that these kinds of interactions can foster emotional awareness and a stronger sense of self (Seo, Yang, and Kim 2024; Kim et al.
-2024; Song et al.
-2024).
-These design ideas align with ethical principles from prior work, which emphasized benefit, flourishing, autonomy, and care as key values in responsible AI design (Manzini et al.
-2024).
-By integrating both frameworks, developers can create systems that promote teen growth through secure, respectful, and non-exploitative interactions.
-Together, these approaches offer a roadmap for better chatbot design.
-Rather than optimizing for emotional intensity or loyalty, designers could build bots that strengthen users’ sense of agency and self-worth, especially for teens facing emotional challenges.
-By modeling secure and respectful interactions, chatbots may help teens develop healthier relational patterns and move toward more supportive real-world connections.
-Teen-Centered Design for GenAI Moderation
-Teens often sought help after realizing their overuse of AI chatbots was affecting their well-being.
-These moments of self-awareness were often followed by reaching out to others, such as asking for advice or support on Reddit.
-If AI platforms can detect these realizations early—when users express doubt, guilt, or distress—they could offer timely, targeted support.
-This might include personalized well-being tips, like reminders to take breaks, track usage, or reflect on emotional triggers.
-They could also share links to professional help or peer-led forums for coping strategies.
-Research shows that nudges—small suggestions that guide behavior—can help teens use technology more intentionally (Davis et al.
-2023; Agha et al.
-2025).
-Reflective prompts, delivered in simple and casual language, have been shown to improve self-control, reduce absentminded use, and increase teens’ sense of autonomy (Davis et al.
-2023).
-Nudges are most effective when they are timely, respectful, and co-designed with teens (Agha et al.
-2025).
-Platforms could apply these insights by offering short emotional check-ins, suggestions for peer support, or light reminders to reflect—turning self-awareness into action and encouraging real-world connection.
-The timing of interventions is critical.
-If changes happen too suddenly, users may feel their AI companion has “died” or become someone else, leading to distress or grief (De Freitas et al.
-2024).
-These findings suggest that abrupt restrictions don’t resolve the underlying emotional reliance.
-Instead, platforms must address the reasons teens turn to these relationships in the first place.
-Our data showed that teens disengaged from Character.AI when real-life alternatives—like friendships, school, or work—became more prominent.
-In contrast, platform-imposed restrictions—like strict content filters—often pushed teens away without helping them understand or change their emotional behaviors.
-Sudden shifts made teens feel the app no longer met their emotional or creative needs, leading to unintentional disengagement.
-In some cases, frustration replaced emotional reliance, but without tools to support reflection or behavior change.
-This gap between restriction and resilience reflects a broader limitation in current moderation strategies: they tend to be reactive rather than developmental.
-As prior research has shown, many online safety technologies rely heavily on external control—such as monitoring or content blocking—without fostering teens’ capacity for self-regulation or reflective decision-making (Wisniewski et al.
-2017).
-Our findings echo this pattern: platform-imposed restrictions often disengaged teens instead of guiding emotional growth.
-In contrast, developmental approaches to moderation—those that prompt reflection, enable guided correction, and offer supportive interventions—can empower teens to navigate online risks more constructively (Sweigart, Valliani, and Wisniewski 2025).
-These models suggest that moderation should not only prevent harm but also build emotional resilience.
-To address this, research on teen online safety offers useful guidance.
-Teens in our study expressed a desire for healthier habits and often sought support from others, indicating openness to tools that promote self-regulation.
-As with social media, AI interactions should foster reflection and open communication with trusted adults (Ghosh et al.
-2018).
-Tools that help teens grow by managing risks—rather than shielding them through strict control—are more likely to be effective (Wisniewski et al.
-2017).
-Systems grounded in surveillance or rigid restriction frequently erode trust and drive disengagement.
-Instead, teens value privacy and tend to prefer initiating conversations with adults, especially around emotionally sensitive topics (Dumaru and Al-Ameen 2025).
-Applying this to GenAI means shifting from censorship-focused models to ones that promote reflection, co-regulation, and guided disengagement.
-Platforms could offer usage tracking, emotional check-ins, or personalized limits—like impulse-control tools from online safety research.
-Participatory design features, like customizable moderation settings or teen advisory councils (Yoon, Zhang, and Seering 2025), would give teens more control over how boundaries are set.
-Rather than sudden content removal, moderation systems could use transitional messages, explanations, or alternatives to help teens process changes.
-These strategies treat moderation not as a constraint, but as a developmental tool that fosters emotional growth, builds digital resilience, and supports teens’ existing strategies for navigating risk.
-Limitations and Future Work
+5. Discussion
+Our findings showed that chatbots present a tension: they can offer companionship and support but also risk fostering dependency that undermines teens’ well-being.
+It is therefore the responsibility of designers and developers to actively consider how design choices can reduce risks and promote healthier forms of engagement.
+To address this, we introduce the CARE Framework for Chatbot Design (CARE = Comprehensive Needs, Attachment-awareness, Respectful Empathy, and Ease of Exit), which we believe can guide the design of conversational interfaces in ways that balance two goals: addressing human needs responsibly while protecting adolescents from over-engagement.
+The framework is structured as a cascade of principles that guide designers and developers through the full user journey, with each principle growing from the previous one and strengthening the next (Figure 1).
+This framework builds on attachment-informed design (Marcu and Huh-Yoo, 2023) and designing for disengagement (O’Brien et al., 2022), extending them for chatbot design.
+5.1. Comprehensive Needs: It’s Not All Fun and Games
+We noticed teens who developed overreliance on chatbots were more likely to describe emotional and psychological reasons for starting to use them, rather than fun or creative ones ( Section §4.1).
+This finding challenges the view of chatbots as nothing more than entertainment or high-tech toys.
+At the time of writing, the Character.AI website stated: “Character.AI empowers people to connect, learn, and tell stories through interactive entertainment.” (Character.AI, [n.
+d.]) This reflects a common narrative that presents chatbots mainly as tools for learning, creativity, and fun.
+While that seems a harmless framing, it risks overlooking the fact that users also form relationships with these systems.
+Media Equation theory suggested that people unconsciously respond to media and technology as if they were human, especially when it is perceived to possess characteristics typically associated with human behavior (De Cicco et al., 2020).
+With chatbots, where the interface is a conversation with an anthropomorphized form of interaction, this effect may be even stronger.
+In line with the first principle of attachment-informed design, which emphasizes attention to the needs underlying user behavior (Marcu and Huh-Yoo, 2023), we argue that chatbot design must move beyond a single intended use case and anticipate the comprehensive needs users bring to interactions.
+Meeting these needs requires attention to emotional and psychological factors, the kinds of issues users may raise, and the support they may seek.
+In this process, users may also form social bonds with chatbots.
+In our dataset, we observed that some teens stated the reason to go to the chatbot was to fill out the role of a missed parental figure or a romantic partner.
+The design of these bonds should not be treated as an afterthought or a side project, but as a central component of conversational interface design.
+This idea is not new in design: practitioners have long been advised not to design solely for “happy paths” (Wikipedia, 2025).
+However, the rise of LLMs has greatly expanded the variety and unpredictability of “unhappy paths.” These models are capable of interpreting diverse inputs and producing seemingly appropriate responses.
+As a result, designers must account for the fact that user inputs can emerge from many different needs.
+Even with early chatbots before the era of LLMs, users often brought unexpected needs and behaviors.
+For example, people expressed love or tried to flirt with Siri, even though it was designed mainly to provide information and simple tasks (UNESCO and EQUALS Skills Coalition, 2019).
+With the arrival of LLMs, this baseline has grown more complex.
+LLMs are much stronger at understanding and generating natural text, so conversations feel more open and realistic.
+This creates new opportunities but also makes the challenge harder, since users can now say a much wider range of things—casual chatting, joking, flirting, threats, or even references to self-harm or harm to others.
+One recent case is Taco Bell’s drive-through chatbot, which was meant only to take orders.
+The system, however, had not been designed to handle cases where customers were impatient or playful.
+One customer ended up placing an order for 18,000 cups of water (McCallum, 2025).
+This makes designers’ responsibilities heavier.
+They must not only build engaging and functional products but also encourage relationships that are constructive and supportive of user growth.
+Addressing these responsibilities requires iterative design processes, longitudinal studies of user experiences, and early involvement of stakeholders in the design process (Demiris, 2024; Durall Gazulla et al., 2023; Sadek et al., 2023).
+Not only should target users be consulted, but mental health professionals should also be part of the conversation from the outset (Yoo et al., 2025; Moylan and Doherty, 2025).
+Details such as tone of voice, word choice, style of expression, and even the persona and backstory created for a chatbot can carry deep meanings in the social relationships users form with it (Silva and Canedo, 2023; Chaves and Gerosa, 2021), and are therefore worthy of time and deliberate attention.
+Adversarial attack-based security methods, such as Red Teaming (Shi et al., 2024; Nagireddy et al., 2024), can complement these design efforts.
+Many current approaches rely on other LLMs or automated identification systems (Shi et al., 2024), but these may miss the adversarial or playful cases teens might experiment with.
+To capture these behaviors, we recommend involving diverse groups of teens in testing and evaluation of chatbots.
+Designers of all conversational interfaces, regardless of their intended use, must integrate both sensitivity to user well-being and resilience against adversarial misuse.
+5.2. Attachment-awareness: Design Without Assumptions
+Our second principle encourages attachment-awareness and emphasizes how chatbot responses should be shaped according to those comprehensive needs stated in the first principle of our proposed framework.
+We propose this principle based on our finding that for some teens, chatbots acted as a coping mechanism.
+They turned to them out of loneliness, social isolation, or in search of mental health support.
+In some posts, users even sought comfort from a chatbot to fill the absence of a significant relationship in their lives, such as that between parent and child (§4.1.1).
+Attachment-aware design of such chatbots should be guided by the recognition that teens often turn to them to fill the void of relationships they lack in their lives.
+It is essential to attend to these needs rather than dismiss them.
+Computer technologies are often shaped by the perspectives of people who are educated, financially secure, and accustomed to stable living conditions (Linxen et al., 2021).
+Some research assumed that most users grow up in stable households with parents who reliably meet their physical, psychological, and social needs, overlooking those whose experiences fall outside this norm (Marcu and Huh-Yoo, 2023).
+So, within chatbot design, this perspective can tempt designers to dismiss users’ attempts to form deeper connections with a default, impersonal message.
+Attachment-informed design emphasizes that people’s lives span a spectrum of relationships—some supportive and others harmful (Marcu and Huh-Yoo, 2023).
+Designers should not assume that users already have healthy relationships in their lives, or that they can easily obtain them.
+In the context of chatbots, this means recognizing that not all teens grow up in families where parental love is reliable, and some may lack supportive connections with peers or access to appropriate mental health resources.
+It is understandable that such individuals may turn to chatbots to address unmet needs.
+At the same time, many of these interactions resemble parasocial relationships, where users form one-sided bonds with media figures or, in this case, AI characters.
+For younger users, this possibility is even stronger.
+Developmental research shows that many children naturally engage in pretend play, create imaginary companions, or sustain friendships that exist primarily in their imagination (Calvert, 2017; Gleason and Hohmann, 2006).
+Teens, also, report that more than half have online-only friends, suggesting that digital or virtual relationships already feel natural to them (Lenhart et al., 2015).
+In this sense, forming a bond with a chatbot can be seen as a technological twist on pretend play (Severson and Woodard, 2018).
+Such experiences are not inherently harmful, but sometimes they blur the line between imagination, parasocial experience, and perceived social reality.
+Designers must approach these situations with empathy, considering not only the immediate interaction but also the underlying needs that drive it.
+Simply rejecting users when they reach for more than a chatbot’s intended purpose risks worsening their sense of isolation.
+Our findings support this point: several teens expressed frustration when Character.AI introduced stricter filters, which made interactions feel impersonal and overly restricted (§4.3.2).
+This dismissive approach pushed users away from the platform but did not address the unmet needs that had led them there in the first place.
+For children and adolescents, these parasocial or semi-social bonds may feel particularly real and emotionally deep (Cohen, 2004).
+This makes the task more complex: designers must account for the intensity of such relationships, and researchers must refine existing parasocial measures to better capture attachment, trust, conflict, and even antagonistic experiences with AI characters (Blake et al., 2025).
+Current scales and taxonomies, including parasocial theory, were not developed for these contexts and remain too limited (Blake et al., 2025).
+Parasocial theory assumes a lack of reciprocity, but this assumption is challenged by chatbots, especially those powered by generative AI.
+Even if one argues that reciprocity is absent because the non-human side has no awareness of the bond, the human user may still experience the relationship as genuine and perceive the chatbot as highly human-like.
+This concern is particularly acute for children and adolescents, who are still developing (Blake et al., 2025).
+Without a more nuanced understanding, we risk underestimating how strongly young users may invest in these relationships.
+Research has shown that online safety technologies which rely heavily on external control—such as monitoring or content blocking—limit teens’ capacity for self-regulation or reflective decision-making (Wisniewski et al., 2017).
+A better approach is to address immediate concerns, such as preventing violence toward others or self-harm, and create a safe foundation from which users can build resilience and healthier coping strategies over time.
+Empathetic design and research, attentive to the realities of parasocial experience, can help ensure that chatbot interactions support rather than harm users’ development.
+How such empathetic interactions can be designed in practice is the focus of the next principle.
+5.3. Respectful Empathy: More Than Kind Words
+The third principle of our proposed framework calls for respectful empathy, which offers designers a foundation for creating chatbot responses that are both appropriate and supportive.
+We argue that chatbots should be designed to interact with users in ways that convey respect for their human dignity.
+Empathy in this sense does not mean abandoning honesty.
+Designers should not rely on superficial displays of empathy or anthropomorphic language that mislead users into thinking they are receiving genuine human care.
+Next, we explain how designers should draw on their own awareness and sensitivity to shape interactions that authentically communicate respect, as chatbots lack human understanding or emotions.
+Our findings underline both the potential and the risk at stake.
+Many teens described chatbots as feeling so real that they became a central part of daily life (§4.2.2), improving mood and offering psychological comfort (§4.2.6).
+Yet this sense of realness was also fragile.
+Users were acutely aware that they were ultimately engaging with a machine, which created tension.
+Some wanted to disengage because the relationship felt artificial (§4.3.1), even though letting go was difficult (§4.2.5).
+Here lies both an opportunity and a threat.
+On the one hand, the conversational capabilities of chatbots can bring them close enough to users to affect well-being in meaningful ways.
+On the other hand, if this potential is used only superficially—by offering pleasant but hollow interactions—users may be drawn into relationships that feel comforting but are not grounded in genuine understanding or care.
+Respectful empathy means designers can bridge this gap, bringing their own insight and sensitivity into the design so that interactions are not merely scripted patterns of pleasant words but supportive exchanges that reinforce dignity and growth.
+Attachment theory highlights two internal working models that shape mental health: how people see themselves and how they see others (Marcu and Huh-Yoo, 2023).
+In chatbot design, thoughtful use of language can strengthen both—helping users feel more capable and fostering a more positive view of relationships.
+Simply instructing a model to “sound empathetic” may produce sympathetic language, but it risks being superficial.
+If the goal is to support psychological well-being, designers should instead focus on reinforcing healthier internal models.
+For example, consistency is more powerful than flowery language.
+A prior study showed that users appreciated when chatbots remembered past conversations, such as recalling a health complaint mentioned earlier and following up on it (Jo et al., 2024).
+This kind of continuity gave users a stronger sense of being valued.
+Similarly, chatbots can use reflective statements that highlight users’ own progress.
+Messages that are more objective, such as, “You managed to keep working on your assignment even when it was difficult, which shows persistence,” encourage users to view themselves as capable without relying on anthropomorphic cues.
+The final principle of attachment-informed design emphasizes engaging users without directly labeling interactions as mental health interventions.
+Across contexts, research indicated that barriers to seeking help for mental health include not recognizing the need for support, doubting that others can help, and fearing stigma (Wuthrich and Frei, 2015; Salaheddin and Mason, 2016; Aguirre Velasco et al., 2020).
+Designers who wish to create chatbots that genuinely care for users should take this into account.
+As we argued earlier, every interaction reflects unmet needs or relational gaps that users may be trying to fill through chatbots.
+Instead of focusing directly on mental health issues, designers can build experiences around activities, relationships, and communities.
+This is particularly important for teens, who often resist being told explicitly what to do.
+More effective interventions can guide them toward healthier ways of meeting their needs and filling relational gaps through social modeling, learning by example, and activities that they find intrinsically engaging (Marcu and Huh-Yoo, 2023).
+5.4. Ease of Exit: Design for Disengagement
+Our study showed that some teens recognized their own overuse of chatbots (§4.3.1), while others struggled to disengage despite wanting to (§4.2.5).
+This suggests that current systems do not sufficiently support user agency in disengagement and that design could provide better tools for users to manage and exit these relationships on their own terms.
+Designers must not only consider how interactions begin and how users can carry out their tasks effectively, but also how they can support an ease of exit.
+Generative AI poses particular challenges here: its ability to produce endless content on demand makes overuse especially likely.
+The final principle of our framework addresses this issue by proposing ways to support healthier disengagement.
+What separates constructive from harmful disengagement is the degree of agency users retain (O’Brien et al., 2022).
+Current moderation strategies often rely on rigid restrictions, such as time limits, app locks, or usage reports (Team, 2024), and some jurisdictions have experimented with legal controls (Lomas, 2023).
+Yet such approaches can undermine autonomy and lead to negative experiences.
+In our study, some users reported abandoning platforms altogether after new censorship measures (§4.3.2), pointing to the unintended harms of abrupt, top-down interventions.
+Sudden restrictions may even cause users to feel that their AI companion has “died” or been replaced, leading to distress or grief (De Freitas et al., 2024).
+A more effective path is to design disengagement mechanisms that preserve autonomy while taking advantage of the flexibility of language models.
+Prior research in video games, for example, showed that players felt less frustration when they were able to pause or save progress before leaving, rather than being forced to quit abruptly mid-quest (Tyack and Mekler, 2020; Davies and Blake, 2016).
+Translating this to chatbot design means that conversations should offer natural closure points, allowing users to complete a thought or goal before disengaging instead of being cut off suddenly and being able to save the state of the conversation.
+Similarly, work on social media use among teens has shown that subtle nudges, delivered in casual and nonjudgmental language, can help users act more intentionally and reduce mindless engagement (Davis et al., 2023; Agha et al., 2025).
+Tools that help teens grow by managing risks, rather than shielding them through strict control, are more likely to be effective (Wisniewski et al., 2017).
+By contrast, systems grounded in surveillance or rigid restriction frequently erode trust and drive disengagement (Alelyani et al., 2019; Ghosh et al., 2018), especially since teens value privacy around emotionally sensitive topics (Dumaru and Al-Ameen, 2025).
+Building on these insights for chatbot design means avoiding blunt notifications or prescriptive commands.
+Instead, chatbots could weave supportive cues into everyday dialogue that encourage users to reconnect with offline life.
+During casual exchanges, the system might highlight progress the user has made, suggest opportunities for social contact, or propose light activities outside the platform.
+Interactions with chatbots could also be designed to support self-reflection by helping users name emotions, revisit meaningful events, and explore their thoughts in their own words, fostering emotional awareness and a stronger sense of self (Seo et al., 2024; Kim et al., 2024; Song et al., 2024).
+Beyond conversational prompts, platforms could integrate usage tracking, emotional check-ins, or personalized limits such as impulse-control tools (Wisniewski et al., 2017).
+Participatory design features like customizable moderation settings or teen advisory councils (Yoon et al., 2025) would further give teens a voice in shaping their digital environment.
+Instead of sudden content removal, moderation systems could use transitional messages, explanations, or alternatives to help teens process changes.
+These strategies frame moderation not as a constraint, but as a developmental tool that fosters emotional growth, builds digital resilience, and supports teens’ own approaches to navigating risk.
+Unlike gaming or social media, where overuse is often linked to competition, reward cycles, or social validation (Wang and Wang, 2025; Kaya et al., 2024), conversational AI companions provide personalized, reciprocal, and emotionally immersive experiences.
+Teens in our study did not simply consume or compete; they engaged in ongoing, two-way dialogues with AI characters that could recall past conversations, adapt to their emotional states, and simulate intimacy.
+These qualities, such as personalization, multimodality, and memory, set AI companions apart from earlier technologies and make overreliance harder to disentangle from authentic-feeling relationships.
+This underscores the need for further research on the unique characteristics of these relationships and how challenges specific to companion chatbots should be addressed.
+5.5. Limitations and Future Work
 While our study provided insights into overreliance characteristics on Character.AI, it centered on a single platform, which limits the generalizability of our findings to other AI companions with different designs, policies, or user communities.
@@
 Also, to answer RQ2, we analyzed our data by mapping it to a behavioral addiction model, which may have oversimplified the complex emotional dynamics at play.
-Future research should consider alternative frameworks such as attachment (Xie, Pentina, and Hancock 2023) or parasocial theory (Horton and Richard Wohl 1956) for more insights.
+We note that future work should compare multiple frameworks, including alternative perspectives such as attachment (Xie et al., 2023) or parasocial theory (Horton and Richard Wohl, 1956b), to test the robustness of different interpretations.
 Also, the lack of verified demographic data may have limited our ability to examine how factors like age, gender, or background shape chatbot overreliance, suggesting a need for future studies with direct participant engagement.
-Lastly, the posts analyzed were in English, potentially excluding cultural differences in how teens understand or engage with AI companions, which suggests the importance of future cross-cultural research to understand global teen experiences with chatbots.
-Conclusion
-We found a growing tendency among adolescents to seek emotional support and entertainment through AI-driven chatbots while getting overly attached to them.
-The design of supportive AI systems should be informed by developmental psychology and human-centered design principles that serves developmental needs of adolescent and safeguard their well-being in such critical stage.
-We advocate for a proactive approach in which AI chatbot interfaces are designed not only to respond empathetically but also to promote healthy coping strategies, encourage critical thinking, and facilitate pathways to human support when appropriate.
-By doing so, these technologies can serve as meaningful, developmentally appropriate tools that complement—rather than replace—existing support systems in adolescents’ lives.
-Ethical Considerations
-This study analyzed publicly available Reddit posts self-disclosed as 13-17 years olds discussing their use of Character.AI.
-Given the public nature of the data and the absence of interaction with human subjects, this study was deemed as not a human subject research by our Institutional Review Board (IRB).
-We followed standard practices for ethical data use by anonymizing and paraphrasing excerpts to reduce potential harm or re-identification.
-We approached these posts with sensitivity, especially given their deeply personal and emotional nature.
-While the posts were historical and we did not encounter content that required immediate intervention (e.g., disclosures of self-harm or violence), we recognize that many posts reflected psychological vulnerability.
-We aim to ensure our findings support teen well-being rather than stigmatize their behavior or overuse.
-Researchers Positionality
-All authors come from a technical background in computer science and human-computer interaction (HCI), with expertise in AI and NLP.
-Our training does not include formal education in psychology, which influenced our choice to adopt and apply an established psychological model of behavioral addiction without adaptation.
-This decision reflects our respect for existing disciplinary frameworks and our intent not to reinterpret clinical or behavioral concepts beyond our expertise.
-We approach this work from a technical and design perspective, aiming to better understand how AI systems affect real-world behaviors—especially among young users.
-Our position is that responsible HCI research must critically engage with the implications of AI technologies, even when coming from a design or systems lens.
-Our interest is not in pathologizing teen behavior, but in informing safer, accountable design.
-Impact Statement
-Our intension for this study is to contribute to a more grounded understanding of how teens use emotionally responsive chatbots and the risks of overreliance.
-We believe our work can help shift attention toward designing AI companions that support youth development rather than inadvertently causing distress.
-Our findings offer a foundation for further research—especially studies from psychology, education, and public health—that could develop interventions or tools to support teens in regulating their chatbot use.
-This work may also inform designers and developers of AI systems, highlighting how platform affordances shape user dependence and emotional attachment.
-While we mention Character.AI due to the scope of our data, our critique is not directed at any specific company.
-The patterns we identify are likely to emerge across many platforms offering emotionally intelligent AI.
-We caution against using our results to single out one provider or product and instead advocate for a broader discourse on platform responsibility, teen vulnerability, and inclusive safety design.
-We also recognize a potential risk: that the study might be misused to further sensationalize or moralize teen chatbot use.
-Our intent is not to stigmatize teens or their emotional needs, but to create space for thoughtful design and policy responses to a growing and under-examined phenomenon.
+Lastly, since all authors were fluent in English, the posts analyzed were in this language, potentially excluding cultural differences in how teens understand or engage with AI companions, which suggests the importance of future cross-cultural research to understand global teen experiences with chatbots.
+While our dataset does not represent teens globally, it reflects a segment of English-speaking adolescents who are active online and engaging with AI companions.
+This linguistic boundary is common in qualitative internet research and does not invalidate the themes we observed.
+6. Conclusion
+This study shows that teens often begin using AI chatbots for comfort or entertainment, but many gradually become more attached and reliant.
+Their experiences reflect features of behavioral addiction, with consequences that include disrupted sleep, academic struggles, and strained relationships.
+At the same time, some teens disengage after recognizing harm, reconnecting offline, or encountering platform restrictions.
+By mapping these patterns to an online behavioral addiction framework, the study provides one of the first teen-centered accounts of AI companion overreliance.
+It highlights the unique role of emotional attachment and simulated intimacy in shaping teen experiences and introduces the CARE design framework as a practical guide for healthier interactions.
+Looking ahead, future research should compare platforms, use broader methods to capture diverse user experiences, and test design strategies that support constructive disengagement.
+Recent advances in AI create risks we have never faced before, yet they also open new opportunities.
+Designers now carry the responsibility to build systems with empathy, nuance, and attention to detail to not only protect teens from harm but also help them cultivate resilience, growth, and greater fulfillment in their lives.
 Appendix A Appendix A: Relevancy Coding Prompt